### 03 Apr 2026 v3

# A Damage-Driven Model for Duchenne Muscular Dystrophy: Early-Stage Dynamics and Invasion Thresholds

Abstract
We introduce a spatially extended mathematical model for Duchenne muscular dystrophy based on a damage-driven paradigm, in which immune recruitment is triggered by tissue injury. The model is formulated as a reaction–diffusion–chemotaxis system describing the interaction between healthy tissue, damaged fibers, immune cells and inflammatory signals. We establish the global well-posedness of the system and investigate the early-stage dynamics through linearization around the healthy equilibrium. Our analysis shows that diffusion does not induce Turing instabilities, so that spatial heterogeneity cannot arise from diffusion-driven mechanisms. Instead, disease progression occurs through invasion processes. We derive explicit conditions for the onset of invasion, interpreted as an effective damage reproduction threshold and characterize the minimal propagation speed of pathological fronts, showing that the dynamics is governed by a pulled-front mechanism. Numerical simulations support the analytical results and confirm the transition between decay and invasion. These results provide a mathematical framework for early-stage disease progression and indicate that spatial spreading arise from the expansion of localized damage rather than from intrinsic pattern-forming mechanisms.
footnotetext: This is a preprint version of the manuscript
1Department of Mathematics, University of Palermo, Via Archirafi 34, Palermo, 90123, Italy
2Department of Engineering, University of Palermo, Viale delle Scienze, Ed. 8, Palermo, 90128, Italy
KEYWORDS: reaction–diffusion–chemotaxis system; traveling fronts; invasion thresholds; immune–tissue interaction; spatial disease dynamics.
1 Introduction
Duchenne muscular dystrophy (DMD) is a severe X-linked genetic disorder characterized by progressive muscle degeneration and loss of functional tissue [10]. The absence of dystrophin leads to increased membrane fragility, causing recurrent micro-damage of muscle fibers. These lesions trigger inflammatory signaling and immune cell recruitment [29, 30, 36], giving rise to a damage–inflammation feedback loop that plays a central role in disease progression [9].
Experimental and clinical evidence indicates that immune activity in DMD is predominantly reactive: inflammatory responses are initiated by tissue damage and tend to subside when damage is resolved. In the pathological context, however, persistent injury prevents full recovery, leading to repeated cycles of degeneration and incomplete repair, with progressive loss of functional tissue [11, 40].
Inflammatory signaling can arise even before macroscopic degeneration becomes evident. Elevated levels of chemokines and inflammatory mediators have been detected in apparently healthy muscle regions of DMD patients [26] and early expression of chemotactic signals has been reported during the initial stages of muscle injury [16]. These observations suggest that early damage and immune activation may precede visible tissue degeneration, motivating the study of early-stage dynamical mechanisms.
Mathematical modeling has been used to investigate immune–tissue interactions in muscular dystrophies. More generally, spatially structured models have been increasingly employed to study the interplay between growth, spatial heterogeneity and biological interactions in complex systems (see, e.g. [1, 24]). Early approaches, including predator–prey-type systems, describe immune cells as proliferating through direct interaction with healthy tissue [5]. While such models capture certain qualitative features, they rely on assumptions that are not fully consistent with experimental observations in DMD. In particular, immune recruitment is primarily triggered by tissue damage and the associated release of inflammatory signals, rather than by direct interaction with intact muscle fibers. Moreover, inflammatory activity tends to decline when damage is resolved, indicating that immune dynamics are not self-sustained in the absence of injury. These features are difficult to reconcile with classical predator–prey formulations, where immune populations grow through direct consumption of healthy tissue and may persist even in the absence of tissue injury.
A step toward greater biological realism was provided by Jarrah et al. [20], who developed a compartmental ODE model of immune–muscle interactions in the mdx mouse, explicitly incorporating necrotic fibers, macrophage subpopulations and regeneration. In this framework, immune recruitment is triggered by tissue damage rather than direct interaction with healthy fibers. Subsequent models, such as the FRiND model [19], further refined this damage-mediated perspective by including detailed recruitment mechanisms and macrophage plasticity.
Despite these advances, most existing models are formulated as systems of ordinary differential equations and focus primarily on temporal dynamics [18]. As a consequence, they do not address how pathological processes propagate across muscle tissue. Spatial aspects of disease progression, including the emergence and spread of damaged regions, remain largely unexplored from a mathematical perspective, despite imaging evidence revealing heterogeneous patterns of degeneration and progressive invasion of affected regions observed in magnetic resonance studies [31, 35].
Spatial propagation phenomena in reaction–diffusion systems have been extensively studied since the classical works of Fisher and Kolmogorov–Petrovsky–Piskunov [12, 21] and further developed in the context of biological invasions and traveling waves [39, 32]. In this work, we develop a mathematical framework for DMD that combines damage-driven immune dynamics with spatial effects. Inspired by existing ODE-based models, we introduce a reaction system incorporating tissue capacity constraints, damage-induced regeneration and saturating immune activation. These mechanisms reflect fundamental biological features of muscle repair and inflammatory regulation [37, 2].
Most existing mathematical models for DMD focus on temporal dynamics and do not explicitly account for spatial propagation of damage. To address this limitation, we develop a spatially extended damage-driven framework that enables the analytical study of invasion thresholds and propagation speeds in the early stage of disease progression.
Within this spatial setting, diffusion accounts for local spreading processes, while directed immune cell migration can be described through chemotactic mechanisms. Such mechanisms have been widely modeled in the context of cell migration and pattern formation, starting from the classical Keller–Segel framework and its extensions [17, 27] and have been incorporated in related models of immune-mediated diseases [23]. Accordingly, we incorporate a chemotactic term describing directed migration of immune cells toward inflammatory signals, reflecting the well-established role of chemokine-guided recruitment in damaged muscle tissue.
At the same time, we emphasize that the present work focuses on the early stage of disease progression, where damage and inflammatory activity remain small. In this regime, chemotactic effects enter only at higher order and do not play a leading role in the onset of invasion. Early spatial propagation is therefore primarily driven by local damage–inflammation feedback mechanisms rather than by directed immune migration. The inclusion of chemotaxis nevertheless provides a structurally consistent framework that can capture more complex spatial behaviors in regimes with stronger inflammation, which will be investigated in future work.
The resulting model is formulated as a reaction–diffusion–chemotaxis system describing the spatial evolution of healthy tissue, damaged fibers, immune cells and inflammatory signals. We first establish basic analytical properties of the system, including positivity, boundedness, and positive invariance of a biologically admissible domain, ensuring the well-posedness and physical consistency of the model.
After nondimensionalization, we characterize the steady states of the system and identify a healthy equilibrium together with parameter-dependent pathological equilibria, when they exist: a healthy state, corresponding to the absence of damage and inflammation, and a diseased state, corresponding to a pathological regime with persistent damage and inflammation. While the healthy equilibrium admits an explicit expression, the diseased equilibrium is determined implicitly through a nonlinear algebraic condition, reflecting the complexity of the underlying feedback mechanisms. While the healthy equilibrium admits an explicit expression, the diseased equilibrium is determined implicitly through a nonlinear algebraic condition. A partial analytical characterization can be obtained in a parameter regime consistent with the early-stage scenario, while a complete analysis remains challenging due to the nonlinear coupling of the system.
The early-stage dynamics of disease progression are investigated through linearization around the healthy equilibrium. We show that diffusion does not induce pattern-forming instabilities of Turing type, indicating that spatial heterogeneity cannot arise through diffusion-driven mechanisms. Instead, disease progression occurs through invasion processes, in which localized damage spreads across the tissue.
We derive analytical conditions for the onset of invasion and characterize the minimal propagation speed of pathological fronts. These theoretical predictions are further supported by numerical simulations of the full nonlinear system. Our results provide a quantitative description of early disease spreading and clarify the mechanisms governing spatial progression in DMD. In particular, we identify an invasion threshold determining whether localized damage can grow from small perturbations of the healthy state and characterize the minimal propagation speed of the resulting pathological fronts. Although the reaction dynamics may exhibit multiple equilibria in certain parameter regimes, we show that, in a biologically relevant regime consistent with early-stage dynamics, the pathological equilibrium can be characterized more precisely. The onset of spatial spreading considered here is governed by a pulled invasion mechanism, controlled by the linearization around the healthy equilibrium.
Organization of the paper.
The paper is organized as follows. In section 2, we introduce the dimensional and nondimensionalized models and discuss their biological interpretation. Section 3 is devoted to the analysis of the spatially homogeneous system, where we establish positivity, invariant regions and characterize the equilibria together with the stability of the healthy state. In section 4, we study the well-posedness of the full reaction–diffusion–chemotaxis system, proving global existence, uniqueness and boundedness of solutions within a biologically admissible domain. Section 5 focuses on the early-stage spatial dynamics. By linearizing the system around the healthy equilibrium, we show the absence of diffusion-driven instabilities and demonstrate that spatial progression occurs through invasion processes. We derive analytical conditions for the invasion threshold and characterize the minimal propagation speed of pathological fronts. In section 6, we present numerical simulations of the full nonlinear system, validating the analytical predictions and illustrating the invasion dynamics in one-dimensional domains. Finally, section 7 discusses the biological implications of our findings and outlines possible directions for future research.
2 The model
In this section, we present the model and discuss its interpretation. We consider four state variables describing the main processes involved in DMD: healthy tissue $H$, damaged tissue $D$, macrophages $M$ and chemokines $C$. This minimal set is chosen to capture the damage-driven and inflammation-mediated nature of the disease, focusing on the feedback between tissue injury and immune response while maintaining analytical tractability.
The model adopts a damage-driven perspective, where immune activation is initiated by tissue injury rather than by direct interaction with healthy fibers. This approach differs from earlier predator–prey-type models, where immune populations grow through direct interaction with intact tissue [5], as well as from more detailed models that distinguish multiple immune subpopulations with different functional roles [19, 20]. In contrast, we adopt an effective description in which regeneration processes and immune activity are represented through aggregated variables, without introducing additional compartments. Biologically, muscle repair involves several intermediate stages, including satellite cell activation, differentiation and fusion, which typically occur on timescales faster than the macroscopic progression of degeneration considered in this work. These processes can therefore be incorporated into an effective regeneration mechanism without introducing additional state variables.
Similarly, the immune response is described through a single macrophage population rather than distinct subtypes. While macrophage polarization plays an important biological role, the dominant effect at the tissue scale is the net inflammatory activity associated with recruited immune cells. Representing this activity through a single effective population captures the leading-order dynamics, while avoiding an unnecessary increase in dimensionality that would complicate the analytical investigation of the system.
We therefore propose the following system:
$\begin{cases}\displaystyle\partial_{\tau}H=D_{H}\Delta H+\left(s+r_{H}D\right)\left(1-\frac{H+D}{K}\right)-\mu_{H}H-a(C)MH,\\[11.00008pt] \displaystyle\partial_{\tau}D=D_{D}\Delta D+a(C)MH-\mu_{D}D-r_{H}D\left(1-\frac{H+D}{K}\right),\\[11.00008pt] \displaystyle\partial_{\tau}M=D_{M}\Delta M-\nabla\!\cdot\!\left(\chi(C)M\nabla C\right)+r_{M}D-\mu_{M}M,\\[11.00008pt] \displaystyle\partial_{\tau}C=D_{C}\Delta C+r_{C}D-\mu_{C}C,\end{cases}$ (2.1)
where:
$a(C)=\bar{a}\dfrac{C+C_{\epsilon}}{k_{c}+C},\qquad\chi(C)=\bar{\chi}\dfrac{C}{k_{\chi}+C}$ (2.2)
and all the parameters are positive. A detailed description of the role of each parameter is given in Table 1.
Parameter Description Biological Role
$D_{H},D_{D}$ Tissue diffusion coefficients Spatial expansion of muscle fibers
$D_{M},D_{C}$ Motility coefficients Diffusion of immune and signaling cells
$s$ Basal growth rate Intrinsic maintenance of healthy tissue
$r_{H}$ Repair rate Conversion rate of damaged fibers back to healthy ones
$K$ Carrying capacity Maximum density of the tissue
$\mu_{H},\mu_{D}$ Natural turnover rates Removal rates of healthy and damaged tissue
$\bar{a}$ Maximum damage rate Peak aggressiveness of macrophages under inflammation
$C_{\epsilon}$ Basal activation Residual damage potential at minimal chemokines levels
$k_{c}$ Half-saturation (damage) Chemokines concentration for half-maximal fiber attack.
$r_{M}$ Macrophage recruitment Rate of immune cell infiltration triggered by damage.
$\mu_{M}$ Macrophage decay Apoptosis or clearance of inflammatory cells.
$\bar{\chi}$ Chemotactic sensitivity Strength of directed macrophage migration toward $C$.
$k_{\chi}$ Chemotactic saturation Sensitivity threshold for the chemical gradient.
$r_{C}$ Chemokines production Secretion rate of signals by damaged tissue.
$\mu_{C}$ Chemokines degradation Apoptosis of chemokines.
Table 1: Description of model parameters and their biological interpretation.
The model incorporates a constraint on the total functional tissue through the factor:
$1-\frac{H+D}{K},$
which represents competition for limited tissue capacity. This does not imply conservation of total tissue mass. In the pathological context of DMD, necrosis, immune-mediated clearance and replacement by adipose or fibrotic tissue lead to an irreversible loss of functional muscle. Accordingly, the carrying capacity parameter represents an effective upper bound on the locally sustainable tissue density rather than a conserved quantity.
The term:
$(s+r_{H}D)\left(1-\frac{H+D}{K}\right)$
describes tissue maintenance and regeneration. The constant contribution $s$ represents basal turnover, while the component proportional to $D$ models damage-induced repair activation, reflecting the biological observation that moderate injury stimulates regenerative mechanisms. The same coefficient $r_{H}$ appears in the damaged tissue equation, ensuring mass transfer consistency between compartments at the effective modeling level.
Immune-mediated damage is represented by the interaction term $a(C)MH$, where the function $a(C)$ describes the dependence of tissue destruction on the inflammatory state. The saturating structure:
$a(C)=\bar{a}\frac{C+C_{\epsilon}}{k_{c}+C}$
captures two key biological features: the presence of a residual inflammatory activity even at low chemokine levels ($C_{\epsilon}$) and the saturation of immune aggressiveness at high signaling concentrations due to physiological limits in cellular activation.
The macrophage equation reflects the damage-driven nature of the immune response. Recruitment is proportional to the amount of damaged tissue ($r_{M}D$), consistent with the release of inflammatory mediators from injured fibers, while linear decay represents clearance and apoptosis. The chemokine equation follows a similar structure, with production driven by tissue damage and degradation accounting for molecular turnover. Together, these equations encode the self-amplifying inflammatory loop characteristic of the disease, in which damage promotes inflammation and inflammation further increases damage.
Spatial effects are introduced through diffusion and chemotaxis. Diffusion terms account for local spreading processes, including migration of immune cells, dispersion of signaling molecules and effective spatial redistribution of tissue components associated with remodeling and expansion. The diffusion coefficients for tissue variables are assumed to be relatively small (see Section 6.1 for details), reflecting the limited motility of muscle fibers compared to immune cells and signaling molecules, while maintaining sufficient regularity for mathematical well-posedness of the system.
Directed migration of macrophages is modeled through the chemotactic flux:
$\nabla\cdot(\chi(C)M\nabla C),$
where the sensitivity function $\chi(C)$ has a saturating form. This structure reflects receptor-mediated sensing mechanisms, with increased responsiveness for weak signals and saturation at high chemokine concentrations due to receptor occupancy limits.
To reduce the number of independent parameters and identify the characteristic scales of the system (2.1), we introduce the following set of dimensionless variables:
$t=\mu_{H}\tau,\quad x=\sqrt{\dfrac{\mu_{H}}{D_{H}}}\xi,\quad h=\dfrac{H}{K},\quad d=\dfrac{D}{K},\quad m=\dfrac{\mu_{M}M}{r_{m}K},\quad c=\dfrac{C}{k_{c}}.$
The temporal scale is chosen according to the characteristic turnover rate of healthy tissue, while the spatial scale is determined by the corresponding diffusion length. Tissue densities are normalized by the carrying capacity and the immune and signaling variables are scaled with respect to their natural production-decay balances.
The resulting system reads as:
$\begin{cases}\displaystyle\partial_{t}h=\Delta h+(\sigma+\rho d)(1-h-d)-h-\alpha\,\dfrac{c+c_{\epsilon}}{1+c}\,mh,\\[10.00002pt] \displaystyle\partial_{t}d=D_{d}\Delta d+\alpha\,\dfrac{c+c_{\epsilon}}{1+c}\,mh-\delta d-\rho d(1-h-d),\\[10.00002pt] \displaystyle\partial_{t}m=D_{m}\Delta m-\chi_{0}\nabla\cdot\!\left(\dfrac{c}{\kappa+c}\,m\nabla c\right)+\nu(d-m),\\[10.00002pt] \displaystyle\partial_{t}c=D_{c}\Delta c+rd-\mu c,\end{cases}$ (2.3)
where:
$D_{d}=\frac{D_{D}}{D_{H}},\qquad D_{m}=\frac{D_{M}}{D_{H}},\qquad D_{c}=\frac{D_{C}}{D_{H}},\qquad\delta=\frac{\mu_{D}}{\mu_{H}},\qquad\nu=\frac{\mu_{M}}{\mu_{H}},\qquad\mu=\frac{\mu_{C}}{\mu_{H}},$
$\sigma=\frac{s}{K\mu_{H}},\qquad\rho=\frac{r_{H}}{\mu_{H}},\qquad r=\frac{r_{C}K}{k_{c}\mu_{H}},\qquad\alpha=\frac{\bar{a}\,r_{M}K}{\mu_{H}\mu_{M}},\qquad\chi_{0}=\frac{\bar{\chi}\,k_{c}}{D_{H}},\qquad\kappa=\frac{k_{\chi}}{k_{c}}.$
The dimensionless parameters represent ratios between competing biological processes and therefore provide insights into the relative importance of the underlying mechanisms. In particular, $\sigma$ describes the strength of basal tissue maintenance relative to the natural turnover rate, while $\rho$ quantifies the efficiency of damage-induced regeneration. The parameter $\alpha$ measures the effectiveness of immune-mediated damage compared to the characteristic tissue timescale, whereas $\delta$ represents the clearance rate of damaged tissue relative to the same scale. The parameters $\nu$ and $\mu$ describe, respectively, the relaxation rates of macrophages and chemokines. Finally, $\chi_{0}$ characterizes the intensity of chemotactic migration relative to diffusion and $\kappa$ represents the chemokine concentration at which chemotactic sensitivity reaches half of its maximal value.
This nondimensional formulation highlights the key feedback mechanisms driving the system while reducing structural complexity, facilitating both analytical investigation and numerical exploration.
3 Local dynamics of the reaction system
In this section, we focus on the reaction terms of (2.3), namely on the corresponding spatially homogeneous dynamical system:
$\begin{cases}\displaystyle\dot{h}=(\sigma+\rho d)(1-h-d)-h-\alpha\,\dfrac{c+c_{\epsilon}}{1+c}\,mh,\\[10.00002pt] \displaystyle\dot{d}=\alpha\,\dfrac{c+c_{\epsilon}}{1+c}\,mh-\delta d-\rho d(1-h-d),\\[10.00002pt] \displaystyle\dot{m}=\nu(d-m),\\[5.0pt] \displaystyle\dot{c}=rd-\mu c.\end{cases}$ (3.1)
3.1 Positive invariance of the physical domain
We first show that the spatially homogeneous system preserves the biologically admissible state space. In particular, healthy and damaged tissue remain nonnegative and cannot exceed the total available tissue fraction, while macrophages and chemokines remain nonnegative.
Theorem 1
Consider system (3.1). Then the set:
$\mathcal{D}=\{(h,d,m,c)\in\mathbb{R}^{4}:\ h\geq 0,\ d\geq 0,\ m\geq 0,\ c\geq 0,\ h+d\leq 1\}$ (3.2)
is positively invariant.
Proof. We verify that the vector field is inward pointing or tangent on each boundary component of $\mathcal{D}$.
(i) Boundary $h=0$. In this case,
$\dot{h}=(\sigma+\rho d)(1-d)\geq 0,$
since $d\in[0,1]$ on $\mathcal{D}$. Hence the vector field points inward or is tangent to the boundary $h=0$.
(ii) Boundary $d=0$. In this case,
$\dot{d}=\alpha\frac{c+c_{\epsilon}}{1+c}mh\geq 0,$
so the vector field points inward or is tangent to the boundary $d=0$.
(iii) Boundary $m=0$. In this case,
$\dot{m}=\nu d\geq 0,$
hence trajectories cannot cross the boundary $m=0$ toward negative values.
(iv) Boundary $c=0$. In this case,
$\dot{c}=rd\geq 0,$
so trajectories cannot cross the boundary $c=0$ toward negative values.
(v) Boundary $h+d=1$. Let $T=h+d$. Summing the first two equations in (3.1) yields:
$\dot{T}=(\sigma+\rho d)(1-T)-h-\delta d.$
Therefore, on the boundary $T=1$,
$\dot{T}|_{T=1}=-h-\delta d\leq 0.$
Thus trajectories cannot cross the boundary $h+d=1$ outward.
Since the vector field points inward or is tangent on every component of the boundary of $\mathcal{D}$, the set $\mathcal{D}$ is positively invariant.
Corollary 2
For any solution with initial condition in $\mathcal{D}$, one has:
$h(t)\geq 0,\quad d(t)\geq 0,\quad m(t)\geq 0,\quad c(t)\geq 0\quad h(t)+d(t)\leq 1,\qquad{\rm for\ all\ }t\geq 0.$
Lemma 3 (Total tissue density is strictly below saturation at equilibrium)
Let $(h^{*},d^{*},m^{*},c^{*})\in\mathcal{D}$ be an equilibrium of system (3.1). Then:
$h^{*}+d^{*}<1.$
Proof. Let $T=h+d$. Summing the first two equations of (3.1) gives:
$\dot{T}=(\sigma+\rho d)(1-T)-h-\delta d.$
At equilibrium,
$0=(\sigma+\rho d^{*})(1-T^{*})-h^{*}-\delta d^{*},$
where $T^{*}=h^{*}+d^{*}$. Hence:
$(\sigma+\rho d^{*})(1-T^{*})=h^{*}+\delta d^{*}.$
Since $\delta>0$, the identity $h^{*}+\delta d^{*}=0$ implies $h^{*}=d^{*}=0$. But this cannot occur at equilibrium, since in that case one would also have $m^{*}=c^{*}=0$ and therefore $\dot{h}=\sigma\neq 0$. Since $\sigma+\rho d^{*}>0$, it follows that
$1-T^{*}>0,$
that is,
$h^{*}+d^{*}<1.$
3.2 Classification of equilibria
We now characterize the equilibria of system (3.1) and discuss their stability properties.
3.2.1 Healthy equilibrium
The system always admits the healthy equilibrium:
$H^{\star}=(h_{0},0,0,0),\qquad\text{where}\quad h_{0}=\frac{\sigma}{1+\sigma}.$ (3.3)
The Jacobian matrix at $H^{\star}$ is:
$J(H^{\star})=\begin{pmatrix}-(1+\sigma)&\dfrac{\rho}{1+\sigma}-\sigma&-\alpha c_{\epsilon}h_{0}&0\\[3.99994pt] 0&-\delta-\dfrac{\rho}{1+\sigma}&\alpha c_{\epsilon}h_{0}&0\\[3.99994pt] 0&\nu&-\nu&0\\ 0&r&0&-\mu\end{pmatrix}$.
The eigenvalues associated with the $h$- and $c$-components are $(1+\sigma)$ and $\mu$, and are therefore strictly negative. The stability of $H^{\star}$ is thus determined by the coupled $(d,m)$ subsystem.
The corresponding characteristic polynomial is:
$\lambda^{2}+a_{1}\lambda+a_{0},$
with:
$a_{1}=\delta+\nu+\frac{\rho}{1+\sigma}>0,\qquad a_{0}=\nu\left(\delta+\frac{\rho}{1+\sigma}-\alpha c_{\epsilon}\frac{\sigma}{1+\sigma}\right).$
Therefore, the sign of $a_{0}$ determines the stability of $H^{\star}$. Introducing the quantity:
$\Theta:=\delta(1+\sigma)+\rho-\alpha c_{\epsilon}\sigma,$ (3.4)
we obtain:
$a_{0}=\frac{\nu}{1+\sigma}\,\Theta.$
The healthy equilibrium $H^{\star}$ is thus locally asymptotically stable if $\Theta>0$ and loses stability at $\Theta=0$.
This threshold coincides with the invasion condition derived from the linearized reaction–diffusion system (see Remark 21).
3.2.2 Pathological equilibria
Besides the healthy equilibrium, solutions with $d^{*}>0$ may arise. At equilibrium, the last two equations of (3.1) yield:
$m^{*}=d^{*},\qquad c^{*}=\frac{r}{\mu}d^{*}.$ (3.5)
Moreover, imposing $\dot{h}+\dot{d}=0$ gives:
$(\sigma+\rho d^{*})(1-h^{*}-d^{*})-h^{*}-\delta d^{*}=0,$
from which we obtain:
$h^{*}=\frac{\sigma-(\sigma+\delta)d^{*}}{1+\sigma}.$ (3.6)
Therefore, any pathological equilibrium must be of the form:
$D^{\star}=\left(\frac{\sigma-(\sigma+\delta)d^{*}}{1+\sigma},\,d^{*},\,d^{*},\,\frac{r}{\mu}d^{*}\right),$ (3.7)
where $d^{*}$ is determined by substituting (3.6) into the condition $\dot{d}=0$. This leads to the quadratic equation:
$G(d^{*})=0,\qquad\qquad{\rm with}\qquad G(d):=g_{2}d^{2}+g_{1}d+g_{0},$ (3.8)
and
$g_{2}$ $=r\big(\rho(1-\delta)-\alpha(\delta+\sigma)\big),$ (3.9)
$g_{1}$ $=-\delta(r+\mu\rho+r\sigma)+\mu\rho-r(\rho-\alpha\sigma)-\alpha\mu(\delta+\sigma)c_{\epsilon},$ (3.10)
$g_{0}$ $=-\mu(\delta+\rho+\delta\sigma)+\alpha\mu c_{\epsilon}\sigma.$ (3.11)
The number of pathological equilibria is determined by the roots of (3.8) for which the corresponding equilibrium $D^{\star}$ lies in the invariant set $\mathcal{D}$ defined in (3.2). Depending on the parameter values, up to two such equilibria may occur. Although a complete analytical classification of (3.8) is difficult in general, a sharper result can be obtained in the regime $\delta\geq 1$, which is consistent with the early-stage scenario considered in this work.
Proposition 4
Assume $\delta\geq 1$ and $\Theta<0$, with $\Theta$ given in (3.4). Then equation (3.8) admits a unique positive root $d^{*}$ and this root satisfies:
$0<d^{*}<\frac{\sigma}{\sigma+\delta}.$ (3.12)
In particular, the corresponding equilibrium $D^{\star}$, as defined in (3.7), belongs to the invariant set $\mathcal{D}$ in (3.2).
Proof. From the definition of $\Theta$ in (3.4), the condition $\Theta<0$ implies $g_{0}>0$, where $g_{0}$ is given in (3.11).
Moreover, since $\delta\geq 1$, we have $g_{2}<0$, see (3.9). Hence $g_{2}g_{0}<0$ and therefore the quadratic polynomial $G(d)$ admits two distinct real roots of opposite sign. In particular, there exists a unique positive root $d^{*}$.
Let:
$\tilde{d}:=\frac{\sigma}{\sigma+\delta}.$
[…]
Using (3.6), this implies $h^{*}>0$. Moreover, since $\delta\geq 1$ and $d^{*}>0$, we have:
$h^{*}+d^{*}=\frac{\sigma+(1-\delta)d^{*}}{1+\sigma}<1.$
Finally, from (3.5) we obtain $m^{*}>0$ and $c^{*}>0$. Hence $D^{\star}\in\mathcal{D}$.
The overall structure of pathological equilibria in the parameter plane $(\delta,\alpha)$ is illustrated in Fig. 1. While multiple equilibria may arise in the general case, the previous result shows that, for $\delta\geq 1$, a unique biologically admissible equilibrium exists above the invasion threshold.
Figure 1: Numerically computed bifurcation diagram in the $(\delta,\alpha)$ parameter plane. The solid black line corresponds to the threshold condition $\Theta=0$, with $\Theta$ given in (3.4), separating regions with different stability properties of the healthy equilibrium. In the white region, the healthy equilibrium is stable and no biologically admissible pathological equilibrium exists. In the red region, the healthy equilibrium is unstable and a stable pathological equilibrium exists. The green region corresponds to a bistable regime, where the healthy equilibrium coexists with two pathological equilibria, one stable and one unstable. The parameters are chosen as $\sigma=1$, $\rho=0.9$, $c_{\epsilon}=0.1$, $\nu=7$, $k=r=1$ and $\mu=168.$
More precisely, the black line corresponds to the threshold condition $\Theta=0$, with $\Theta$ given in (3.4), which determines the loss of stability of the healthy equilibrium.
In the red region, the healthy equilibrium is stable and no biologically admissible pathological equilibrium exists. In the blue region, the healthy equilibrium is unstable and a stable pathological equilibrium emerges, corresponding to a regime of persistent damage.
For sufficiently small values of $\delta$, a bistable regime appears (green region), in which the healthy equilibrium coexists with two pathological equilibria, one stable and one unstable. This indicates the possibility of threshold-driven transitions between recovery and sustained degeneration.
We emphasize that this regime highlighted in the zoomed region of Fig. 1 corresponds to small values of $\delta$, associated with slow damage clearance and it is therefore not representative of the early-stage dynamics considered in this work.
A complete analytical description of (3.8) remains challenging in general, due to the additional constraints $0<h^{*}+d^{*}\leq 1$ and the nonlinear dependence on the parameters. In particular, a full characterization of the below-threshold regime appears to require additional conditions on the coefficients of (3.8).
For this reason, in the following we focus on the dynamics near the healthy equilibrium and on the early-stage regime, where the onset of disease progression can be analyzed through linear mechanisms.
4 Well-posedness of the reaction-diffusion-chemotaxis system
In this section, we establish the global well-posedness of the reaction–diffusion–chemotaxis system (2.3). In particular, we prove the existence, uniqueness and boundedness of solutions, together with the preservation of a biologically admissible domain, ensuring that all components remain non-negative and that the saturation constraint $h+d\leq 1$ is satisfied for all times. The analysis relies on a Leray–Schauder fixed point argument, combined with a priori estimates and parabolic regularity techniques. These results provide the mathematical framework required for the subsequent investigation of the dynamical properties of the model.
For the well-posedness analysis, we rewrite the dimensionless system in the following form:
$\left\{\begin{array}[]{ll}\displaystyle\partial_{t}h=\Delta h+(\sigma+\rho d)(1-h-d)-h-A(m,c)h\,,&(t,x)\in\Omega_{T}\cr\displaystyle\partial_{t}d=D_{d}\Delta d+A(m,c)h-\delta d-\rho d(1-h-d)\,,&(t,x)\in\Omega_{T}\cr\displaystyle\partial_{t}m=D_{m}\Delta m-\nabla\cdot\left(\chi(m,c)\nabla c\right)+\nu(d-m)\,,&(t,x)\in\Omega_{T}\cr\displaystyle\partial_{t}c=D_{c}\Delta c+rd-\mu c,&(t,x)\in\Omega_{T}\end{array}\right$. (4.1)
with:
$\chi(m,c)=\chi_{0}\dfrac{m\,c}{\kappa+c},\quad A(m,c)=\alpha\dfrac{c+c_{\epsilon}}{1+c}\,m\,$ (4.2)
and we denote by $\Omega_{T}=(0,T)\times\Omega$, where $\Omega$ is a bounded domain in $\mathbb{R}^{n}$ $(n\in\mathbb{N},\ n\geq 1)$ with smooth, specified later, boundary $\partial\Omega$. We write $\mathbf{n}=\mathbf{n}(s)$ for the outward normal to $\partial\Omega$ at a point $s$ of the boundary and impose the following Neumann boundary condition on $\partial\Omega$:
$\begin{cases}\mathbf{n}(s)\cdot\bm{\nabla}h(t,s)=0\,&\qquad(t,s)\in(0,T)\times\partial\Omega\\ \mathbf{n}(s)\cdot\bm{\nabla}d(t,s)=0\,&\qquad(t,s)\in(0,T)\times\partial\Omega\\ \mathbf{n}(s)\cdot\bm{\nabla}m(t,s)=0\,&\qquad(t,s)\in(0,T)\times\partial\Omega\\ \mathbf{n}(s)\cdot\bm{\nabla}c(t,s)=0\,&\qquad(t,s)\in(0,T)\times\partial\Omega\end{cases}$ (4.3)
Finally, we impose the following non-negative initial conditions:
$\begin{cases}h(t=0,x)=h_{in}(x)\,,&\qquad x\in\Omega\cr d(t=0,x)=d_{in}(x)\,,&\qquad x\in\Omega\cr m(t=0,x)=m_{in}(x)\,,&\qquad x\in\Omega\cr c(t=0,x)=c_{in}(x)\,,&\qquad x\in\Omega\cr\end{cases}$ (4.4)
and we assume that the following condition holds:
$h_{in}(x)+d_{in}(x)\leq 1\,,\quad x\in\Omega\,.$ (4.5)
Moreover, we recall that all the system parameters are positive. We introduce the functional framework used throughout the section (see [15, 22]).
For $1\leq p\leq\infty$, we denote by $L^{p}(\Omega)$ the usual Lebesgue spaces with norm $\|\cdot\|_{L^{p}(\Omega)}$ and by $L^{p}(\Omega_{T})=L^{p}((0,T);L^{p}(\Omega))$ the corresponding space-time spaces, equipped with the norm $\|\cdot\|_{L^{p}(\Omega_{T})}$.
More generally, for $0\leq t_{0}<T$, we set $\Omega_{t_{0},T}=(t_{0},T)\times\Omega$ and define $L^{p}(\Omega_{t_{0},T})$ accordingly.
We write $W^{r,p}(\Omega)$ for the Sobolev spaces of order $r$ with integrability $p$. For the parabolic setting, we consider the space:
$W^{1,2}_{p}(\Omega_{t_{0},T})=\left\{v:\ \partial_{t}^{r}\partial_{x}^{s}v\in L^{p}(\Omega_{t_{0},T}),\quad 2r+s\leq 2\right\},$
endowed with the norm:
$\|v\|_{W^{1,2}_{p}(\Omega_{t_{0},T})}=\sum_{2r+s\leq 2}\|\partial_{t}^{r}\partial_{x}^{s}v\|_{L^{p}(\Omega_{t_{0},T})}.$
We also denote by $C^{k}(\overline{\Omega})$ the space of functions with continuous derivatives up to order $k$ and by $C^{0}(\Omega_{T})$ the space of continuous functions in $\Omega_{T}$.
We finally introduce the space $C^{0,1}=C^{0,1}([0,T]\times\overline{\Omega})$ of functions that are continuous together with their spatial gradients, endowed with the norm:
$\|u\|_{C^{0,1}}=\sup_{t\in[0,T]}\|u(t,\cdot)\|_{L^{\infty}(\Omega)}+\sup_{t\in[0,T]}\|\nabla u(t,\cdot)\|_{L^{\infty}(\Omega)}.$
For a Banach space $X$, we denote by $X]^{m}$ the Cartesian product of $m$ copies of $X$, endowed with the natural norm.
Our main well-posedness result is stated in the following theorem.
Theorem 5
Let $\Omega$ be a smooth ($C^{2+\eta}$, for some $\eta>0$) bounded connected open subset of $\mathbb{R}^{n}$, with $n\in\mathbb{N}$ and $n\geq 1$. For $\tilde{p}>n+2$, let the non-negative initial conditions $(h_{in},d_{in},m_{in},c_{in})\in[W^{(2-2/\tilde{p}),\tilde{p}}(\Omega)]^{4}$, with (4.5), then the system (4.1) admits a unique global strong solution which is non-negative in each component, with $h+d\leq 1$ in $\Omega$ and bounded in time. More precisely, we have that:
$\partial_{t}(h,d,m,c)\,,\,\,\Delta(h,d,m,c)\,\in[L^{\tilde{p}}\left(\Omega_{T}\right)]^{4}\,\,,\,\,\quad\quad\nabla\cdot\left(\chi(m,c)\nabla c\right)\in L^{\tilde{p}}\left(\Omega_{T}\right)\,$
and $(h,d,m,c)\in[C^{0,1}]^{4}$, for all $T>0$.
Furthermore, if $(h_{in},d_{in},m_{in},c_{in})\in[C^{2}(\overline{\Omega})]^{4}$, non negative with (4.5), and $\nabla h_{in}\cdot\mathbf{n}=\nabla d_{in}\cdot\mathbf{n}=\nabla m_{in}\cdot\mathbf{n}=\nabla c_{in}\cdot\mathbf{n}=0$ on $\partial\Omega$, then the solution to the system (4.1) is classical, i.e., $\forall T>0$,
$\partial_{t}(h,d,m,c)\,,\,\Delta(h,d,m,c)\in[C^{0}\left(\Omega_{T}\right)]^{4}\,,\,\,\quad\quad\nabla\cdot\left(\chi(m,c)\nabla c\right)\in C^{0}\left(\Omega_{T}\right)\,.$
Moreover, there exists a constant $\overline{C}>0$, depending on $\Omega,h_{in},d_{in},m_{in},c_{in}$ and on all the system parameters, such that the unique global solution of the system (4.1) satisfies:
$\sup_{t\geq 0}\left\|\left(h(t,\cdot)\,,\,d(t,\cdot)\,,\,m(t,\cdot)\,,\,c(t,\cdot)\right)\right\|_{[W^{1,\infty}(\Omega)]^{4}}\leq\overline{C}\,$ (4.6)
and
$\limsup_{t\rightarrow+\infty}\left\|\left(h(t,\cdot)\,,\,d(t,\cdot)\,,\,m(t,\cdot)\,,\,c(t,\cdot)\right)\right\|_{[W^{1,\infty}(\Omega)]^{4}}\leq\overline{C}\,.$ (4.7)
The proof of Theorem 5 is divided into three parts. In section 4.1 we prove the global existence of solutions by means of a Leray–Schauder fixed point argument. In section 4.2 we establish uniqueness, while in section 4.3 we derive the uniform-in-time bounds (4.6)–(4.7). The argument builds on the approach developed in [6, 13], suitably adapted to the specific nonlinear and coupled structure of the reaction–diffusion system (4.1).
We recall the following embedding lemma, which will be used in the sequel (see [22, 33]).
Lemma 6
Let $1<p<\infty$, $0\leq t_{0}<T$ and $\Omega$ be a bounded domain in $\mathbb{R}^{n}$. Then there exists a constant $C_{T}>0$, depending on $T-t_{0}$, $\Omega$, $p$ and $n$, such that for all $u\in W^{1,2}_{p}(\Omega_{t_{0},T})$ the following estimates hold:
[…]
  • $W^{1,2}_{p}(\Omega_{t_{0},T})$ is compactly embedded in $C^{0,1}$, if $p>n+2$.
4.1 Global existence of solutions
In this section, we prove the global existence of non-negative solutions to system (4.1) by means of a Leray–Schauder fixed point argument. We also derive uniform bounds in the $L^{\infty}$ norm.
The proof is structured as follows. We first construct a suitable map $S$ and establish its compactness and continuity. We then introduce the set $\Phi$ of its fixed points and prove that its elements are non-negative and satisfy the saturation constraint. Finally, we show that $\Phi$ is bounded and conclude the existence result.
4.1.1 Construction of the map $S$
We begin by constructing a suitable solution operator $S$, which will be used to apply the Leray–Schauder fixed point theorem, as follows:
$S:[C^{0,1}]^{4}\times[0,1]\longrightarrow[C^{0,1}]^{4}\,.$ (4.10)
Given $(h,d,m,c)\in[C^{0,1}]^{4}$ and $\lambda\in[0,1]$, we define:
$S(h,d,m,c,\lambda)=(\lambda\,h^{\ast},\lambda\,d^{\ast},\lambda\,m^{\ast},\lambda\,c^{\ast}),$ (4.11)
where $(h^{\ast},d^{\ast},m^{\ast},c^{\ast})$ is obtained by solving, in sequence, the following linear parabolic problems.
For a given $d\in C^{0,1}$, we consider the problem:
$\begin{cases}\partial_{t}c^{\ast}=D_{c}\Delta c^{\ast}-\mu c^{\ast}+rd_{+}&\quad(t,x)\in\Omega_{T}\\ \mathbf{n}\cdot\bm{\nabla}c^{\ast}=0&\quad(t,s)\in(0,T)\times\partial\Omega\\ c^{\ast}(t=0,x)=c_{in}(x)&\quad x\in\Omega\end{cases}$ (4.12)
where $d_{+}=\max(0,d)$. Once $c^{\ast}$ is determined, for given $m,c\in C^{0,1}$ we solve:
$\begin{cases}\partial_{t}m^{\ast}=D_{m}\Delta m^{\ast}-\nabla\cdot\left(\chi(m_{+},c_{+})\nabla c^{\ast}\right)+\nu(d_{+}-m^{\ast})&(t,x)\in\Omega_{T}\\ \mathbf{n}\cdot\bm{\nabla}m^{\ast}=0&(t,s)\in(0,T)\times\partial\Omega\\ m^{\ast}(x,0)=m_{in}(x)&x\in\Omega\end{cases}$ (4.13)
where $m_{+}=\max(0,m)$ and $c_{+}=\max(0,c)$. Next, for given $h,d\in C^{0,1}$, we solve:
$\begin{cases}\partial_{t}d^{\ast}=D_{d}\Delta d^{\ast}-\delta d^{\ast}-\rho d^{\ast}+\rho d_{+}\overline{(h+d)}_{+}+A(m^{\ast},c_{+})h_{+}&(t,x)\in\Omega_{T}\\ \mathbf{n}\cdot\bm{\nabla}d^{\ast}=0&(t,s)\in(0,T)\times\partial\Omega\\ d^{\ast}(x,0)=d_{in}(x)&x\in\Omega\end{cases}$ (4.14)
where $h_{+}=\max(0,h)$ and $\overline{(h+d)}_{+}=\max(0,h+d)\wedge\min(1,h+d)$.
Finally, we consider:
$\begin{cases}\partial_{t}h^{\ast}=\Delta h^{\ast}-h^{\ast}-A(m_{+},c_{+})h^{\ast}+\sigma(1-\overline{(h+d)}_{+})\\ \phantom{aaaa}+\rho d^{\ast}(1-\overline{(h+d)}_{+})&(t,x)\in\Omega_{T}\\ \mathbf{n}\cdot\bm{\nabla}h^{\ast}=0&(t,s)\in(0,T)\times\partial\Omega\\ h^{\ast}(x,0)=h_{in}(x)&x\in\Omega\end{cases}$ (4.15)
By construction, the map $S$ associates to each $(h,d,m,c,\lambda)$ an element of $C^{0,1}]^{4}$.
Moreover, from the definition of $S$ given by (4.12)–(4.15) and the maximal regularity theory for parabolic equations with Neumann boundary conditions (see [6, 22]), we obtain:
$\|(h^{\ast},d^{\ast},m^{\ast},c^{\ast})\|_{[W^{1,2}_{p}(\Omega_{T})]^{4}}\leq K^{\ast},$ (4.16)
where $K^{\ast}$ depends on $T$, $\Omega$, $n$, $p$, the initial data, the parameters of the system and the $C^{0,1}$ norms of $h$, $d$, $m$ and $c$.
Since $p>n+2$, the space $W^{1,2}_{p}(\Omega_{T})$ is compactly embedded in $C^{0,1}$ (Lemma 6) and therefore:
$(h^{\ast},d^{\ast},m^{\ast},c^{\ast})\in[C^{0,1}]^{4}.$
4.1.2 Compactness and continuity of $S$
To apply the Leray–Schauder fixed point theorem, we show that the map $S$ is compact and continuous in $C^{0,1}]^{4}$. This is achieved by establishing precompactness of $S$ on bounded sets and then proving continuity in the following lemmas.
Lemma 7
Under the hypotheses of Theorem 5, the map $S$ maps bounded sets of $C^{0,1}]^{4}\times[0,1]$ into precompact sets of $C^{0,1}]^{4}$.
Proof. The claim follows from (4.16). Since $p>n+2$, the space $W^{1,2}_{p}(\Omega_{T})$ is compactly embedded in $C^{0,1}$ by Lemma 6, which yields the desired precompactness.
Lemma 8
Under the hypotheses of Theorem 5, the map $S$ is continuous.
Proof. Let $(h_{1},d_{1},m_{1},c_{1}),(h_{2},d_{2},m_{2},c_{2})\in[C^{0,1}]^{4}$.
We start by estimating the difference $c^{\ast}_{1}-c^{\ast}_{2}$. From (4.12), using maximal regularity (see [22]) together with the embedding properties of $W^{1,2}_{p}$ (Lemma 6) for $p>n+2$, we obtain:
$\|c_{1}^{\ast}-c_{2}^{\ast}\|_{C^{0,1}}\leq C^{\ast}_{T}\|d_{1+}-d_{2+}\|_{L_{p}(\Omega_{T})}\leq C^{\ast}_{T}\|d_{1}-d_{2}\|_{C^{0,1}}\,$, (4.17)
where $C^{\ast}_{T}$ denotes a constant depending on $T$, $\Omega$, $n$, the initial data and the parameters of the system.
We now turn to the estimate of $m_{1}^{\ast}-m_{2}^{\ast}$, using (4.13). The following pointwise estimate holds:
$\left|\nabla\cdot\left(\chi(m_{1+},c_{1+})\nabla c_{1}^{\ast}-\chi(m_{2+},c_{2+})\nabla c_{2}^{\ast}\right)\right|\leq\\ \leq\chi_{0}|m_{1+}|\,|\Delta(c_{1}^{\ast}-c_{2}^{\ast})|+\chi_{0}\,|\nabla m_{1+}|\,|\nabla c_{1}^{\ast}-\nabla c_{2}^{\ast}|+\\ \qquad+\chi_{0}|m_{1+}-m_{2+}|\,|\Delta c_{2}^{\ast}|+\chi_{0}|\nabla(m_{1+}-m_{2+})|\,|\nabla c_{2}^{\ast}|\,.$
Applying the same arguments used for $c_{1}^{\ast}-c_{2}^{\ast}$, we deduce:
$\|m_{1}^{\ast}-m_{2}^{\ast}\|_{C^{0,1}}\leq C_{T}\left(\|d_{1}-d_{2}\|_{C^{0,1}}+\|m_{1}-m_{2}\|_{C^{0,1}}\right)\,$, (4.18)
where $C^{\ast}_{T}$ depends on $T$, $\Omega$, $n$, the initial data and the parameters of the system.
We finally estimate $d_{1}^{\ast}-d_{2}^{\ast}$ and $h_{1}^{\ast}-h_{2}^{\ast}$ using (4.14) and (4.15). We first establish the following auxiliary estimates:
$\left|A(m^{\ast}_{1},c_{1+})h_{1+}-A(m_{2}^{\ast},c_{2+})h_{2+}\right|\leq\alpha(1+c_{\epsilon})\left(h_{2+}|m_{1}^{\ast}-m_{2}^{\ast}|+|m_{1}^{\ast}||h_{1+}-h_{2+}|\right)+\\ +\alpha\,h_{2+}\,|m_{1}^{\ast}|\,|1-c_{\epsilon}||c_{1+}-c_{2+}|\,$, (4.19)
$\left|\rho d_{1+}\overline{(h_{1}+d_{1})}_{+}-\rho d_{2+}\overline{(h_{2}+d_{2})}_{+}\right|\leq\\ \leq\rho d_{2+}\left(|h_{1+}-h_{2+}|+|d_{1+}-d_{2+}|\right)+\rho(h_{1+}+d_{1+})|d_{1+}-d_{2+}|\,$, (4.20)
$\left|A(m_{1+},c_{1+})h_{1}^{\ast}-A(m_{2+},c_{2+})h_{2}^{\ast}\right|\leq\alpha(1+c_{\epsilon})\left(|h_{2}^{\ast}||m_{1+}-m_{2+}|+m_{1+}|h_{1}^{\ast}-h_{2}^{\ast}|\right)+\\ +\alpha\,m_{1+}\,|h_{2}^{\ast}|\,|1-c_{\epsilon}||c_{1+}-c_{2+}|\,$, (4.21)
$|\overline{(h_{1}+d_{1})}_{+})-\overline{(h_{2}+d_{2})}_{+}|\leq(|h_{1}-h_{2}|+|d_{1}-d_{2}|)\,$, (4.22)
$\left|\rho d_{1}^{\ast}(1-\overline{(h_{1}+d_{1})}_{+})-\rho d_{2}^{\ast}(1-\overline{(h_{2}+d_{2})}_{+})\right|\leq\\ \leq\rho|d_{2}^{\ast}|\left(|h_{1}-h_{2}|+|d_{1}-d_{2}|\right)+\rho(1+h_{1+}+d_{1+}))|d_{1}^{\ast}-d_{2}^{\ast}|\,.$ (4.23)
Using (4.19)–(4.23) together with the same arguments as before, we obtain:
$\|d_{1}^{\ast}-d_{2}^{\ast}\|_{C^{0,1}}\leq C_{T}\left(\|h_{1}-h_{2}\|_{C^{0,1}}+\|d_{1}-d_{2}\|_{C^{0,1}}+\|m_{1}-m_{2}\|_{C^{0,1}}+\|c_{1}-c_{2}\|_{C^{0,1}}\right)\,$, (4.24)
and
$\|h_{1}^{\ast}-h_{2}^{\ast}\|_{C^{0,1}}\leq C_{T}\left(\|h_{1}-h_{2}\|_{C^{0,1}}+\|d_{1}-d_{2}\|_{C^{0,1}}+\|m_{1}-m_{2}\|_{C^{0,1}}+\|c_{1}-c_{2}\|_{C^{0,1}}\right)\,$, (4.25)
where $C^{\ast}_{T}$ depends on $T$, $\Omega$, $n$, the initial data and the parameters of the system.
Combining (4.17), (4.18), (4.24) and (4.25), we conclude that the map $S$ is continuous.
4.1.3 The set $\Phi$
To apply the Leray–Schauder theorem, we introduce the set $\Phi$:
$\Phi=\left\{(h,d,m,c)\in[C^{0,1}]^{4}:S(h,d,m,c,\lambda)=(h,d,m,c),\;0<\lambda\leq 1\right\}.$ (4.26)
By definition, if $(h,d,m,c)\in\Phi$, then $h=\lambda h^{\ast}$, $d=\lambda d^{\ast}$, $m=\lambda m^{\ast}$ and $c=\lambda c^{\ast}$, where $h^{\ast}$, $d^{\ast}$, $m^{\ast}$ and $c^{\ast}$ solve (4.15), (4.14), (4.13) and (4.12), respectively.
Using again the definition of $\Phi$, and multiplying equations (4.15), (4.14), (4.13) and (4.12) by $\lambda$, we deduce that, if $(h,d,m,c)\in\Phi$, then $(h,d,m,c)$ satisfies the following system:
$\begin{cases}\displaystyle\partial_{t}h=\Delta h-h-A(m_{+},c_{+})h+\lambda\sigma(1-\overline{(h+d)}_{+})+\rho d(1-\overline{(h+d)}_{+}),&(t,x)\in\Omega_{T}\\ \displaystyle\partial_{t}d=D_{d}\Delta d-(\delta+\rho)d+A(m,c_{+})h+\rho\lambda d_{+}\overline{(h+d)}_{+},&(t,x)\in\Omega_{T}\\ \displaystyle\partial_{t}m=D_{m}\Delta m-\nabla\cdot\left(\chi(m_{+},c_{+})\nabla c\right)+\nu(\lambda d_{+}-m),&(t,x)\in\Omega_{T}\\ \displaystyle\partial_{t}c=D_{c}\Delta c-\mu c+r\lambda d_{+}\,,&(t,x)\in\Omega_{T}\end{cases}$ (4.27)
with boundary conditions:
$\mathbf{n}\cdot\bm{\nabla}h(t,s)=\mathbf{n}\cdot\bm{\nabla}d(t,s)=\mathbf{n}\cdot\bm{\nabla}m(t,s)=\mathbf{n}\cdot\bm{\nabla}c(t,s)=0\,\,\quad(t,s)\in(0,T)\times\partial\Omega\$(4.28)
and initial conditions:
$\begin{cases}h(0,x)=\lambda h_{in}(x)&x\in\Omega\\ d(0,x)=\lambda d_{in}(x)&x\in\Omega\\ m(0,x)=\lambda m_{in}(x)&x\in\Omega\\ c(0,x)=\lambda c_{in}(x)&x\in\Omega\end{cases}$ (4.29)
which are non-negative and satisfy the saturation condition (4.5). We first establish the following Lemma.
Lemma 9
If $(h,d,m,c)$ are in $\Phi$, then they are non-negative.
Proof. Since $c_{in}\geq 0$ and $d_{+}\geq 0$, then $c\geq 0$. To prove that $m\geq 0$, we multiply the equation for $m$ by $m_{-}^{2}$, with $m_{-}=\max(0,-m)$ and integrate in space and time. Recalling that $m_{in}\geq 0$, we obtain:
$\frac{1}{3}\int_{\Omega}m_{-}^{3}dx=2D_{m}\int_{0}^{t}\int_{\Omega}m_{-}|\nabla m_{-}|^{2}dxdt-\nu\int_{0}^{t}\int_{\Omega}mm_{-}^{2}dxdt\\ +\nu\lambda\int_{0}^{t}\int_{\Omega}d_{+}m_{-}^{2}dxdt-2\chi_{0}\int_{0}^{t}\int_{\Omega}\frac{cm_{-}m_{+}}{1+c}\nabla m_{-}\cdot\nabla cdxdt\\ \geq 2\int_{0}^{t}\int_{\Omega}m_{-}|\nabla m_{-}|^{2}dxdt+\nu\int_{0}^{t}\int_{\Omega}m_{-}^{3}dxdt\geq 0\,$,
and consequently $m_{-}=0$ for each $t\in[0,T]$.
We proceed analogously for $h$ and $d$, obtaining:
$\frac{1}{3}\int_{\Omega}d_{-}^{3}dx=2D_{d}\int_{0}^{t}\int_{\Omega}d_{-}|\nabla d_{-}|^{2}dxdt+(\delta+\rho)\int_{0}^{t}\int_{\Omega}d_{-}d_{-}^{2}+\\ +\lambda\rho\int_{0}^{t}\int_{\Omega}d_{+}\overline{(h+d)}_{+}d_{-}^{2}dxdt+\int_{0}^{t}\int_{\Omega}A(m,c_{+})h_{+}d_{-}^{2}dxdt\geq 0\,$,
$\frac{1}{3}\int_{\Omega}h_{-}^{3}dx=2\int_{0}^{t}\int_{\Omega}h_{-}|\nabla h_{-}|^{2}dxdt+\lambda\sigma\int_{0}^{t}\int_{\Omega}(1-\overline{(h+d)}_{+})h_{-}^{2}dxdt\\ +\int_{0}^{t}\int_{\Omega}h_{-}^{3}dxdt+\int_{0}^{t}\int_{\Omega}A(m_{+},c_{+})h_{-}^{3}dxdt+\rho\int_{0}^{t}\int_{\Omega}d\,(1-\overline{(h+d)}_{+})h_{-}^{2}dxdt\geq 0\,$,
using that $h_{in}\geq 0$ and $d_{in}\geq 0$. Hence $h_{-}=0$ and $d_{-}=0$ for each $t\in[0,T]$.
By Lemma 9, system (4.27) reduces to:
$\begin{cases}\displaystyle\partial_{t}h=\Delta h-h-A(m,c)h+\lambda\sigma(1-\overline{(h+d)}_{+})+\rho d(1-\overline{(h+d)}_{+}),&(t,x)\in\Omega_{T}\\ \displaystyle\partial_{t}d=D_{d}\Delta d-(\delta+\rho)d+A(m,c)h+\rho\lambda d\overline{(h+d)}_{+},&(t,x)\in\Omega_{T}\\ \displaystyle\partial_{t}m=D_{m}\Delta m-\nabla\cdot\left(\chi(m,c)\nabla c\right)+\nu(\lambda d-m),&(t,x)\in\Omega_{T}\\ \displaystyle\partial_{t}c=D_{c}\Delta c-\mu c+r\lambda d\,,&(t,x)\in\Omega_{T}\end{cases}$ (4.30)
with boundary conditions (4.28) and non-negative initial data (4.29) with condition (4.5).
4.1.4 Boundedness of $\Phi$
We now show that $\Phi$ is bounded in $C^{0,1}]^{4}$ (see Lemma 13 below). To this end, we first establish two auxiliary lemmas describing the regularity of $h$, $d$, $m$, and $c$ for $(h,d,m,c)\in\Phi$.
We then derive a sequence of a priori estimates, which will be used to prove the boundedness of the set of fixed points and to conclude the existence argument.
Lemma 10
Let $(h,d,m,c)\in\Phi$. Under the hypotheses of Theorem 5, there exists a constant $H_{1}$, depending on $T$, $\Omega$, $n$, the initial data and the parameters of the system, but not on $\lambda$, such that
$\sup_{t\in[0,T]}\|(h(t,\cdot),d(t,\cdot),m(t,\cdot),c(t,\cdot))\|_{L^{1}(\Omega)}\leq H_{1}\,.$ (4.31)
Proof.
Integrating the equations for $h$ and $d$ in (4.30) and summing, we obtain:
$\partial_{t}\int_{\Omega}(h+d)dx\leq\sigma\int_{\Omega}(1-\overline{(h+d)}_{+})dx\leq\sigma|\Omega|\,.$ (4.32)
Here $|\Omega|$ denotes the measure of the domain. Applying Gronwall’s lemma and using that $h_{in}+d_{in}\leq 1$, we conclude that $h$ and $d$ are bounded in $L^{1}(\Omega)$. Integrating the equations for $m$ and $c$ in (4.30), we obtain:
$\frac{d}{dt}\int_{\Omega}m\,dx\leq-\nu\int_{\Omega}m\,dx+\nu\|d\|_{L^{1}(\Omega)}\,$, (4.33)
and
$\frac{d}{dt}\int_{\Omega}c\,dx\leq-\mu\int_{\Omega}c\,dx+r\|d\|_{L^{1}(\Omega)}\,.$ (4.34)
Applying Gronwall’s lemma, we obtain (4.31).
We now derive a regularity estimate for $c$, which will be used in the sequel.
Lemma 11
Let $(h,d,m,c)\in\Phi$. Under the hypotheses of Theorem 5, there exists a constant $H^{\ast}$, depending on $\Omega$, $T$, $n$, the initial data and the parameters of the system, but not on $\lambda$, such that
$\|\nabla c\|_{L^{(n+2)/(n+1)}(\Omega_{T})}\leq H^{\ast}\,.$ (4.35)
Proof. From the equation for $c$ in (4.30), we may write
$\nabla c(t)=e^{-\mu t}\nabla e^{tD_{c}\Delta}(\lambda c_{in})+r\lambda\int_{0}^{t}e^{-\mu(t-s)}\nabla e^{(t-s)D_{c}\Delta}d(s)\,ds.$
Using classical $L^{p}$-$L^{q}$ estimates (see [14]), we obtain
$\|\nabla c(t)\|_{L^{\frac{n+1}{n+2}}(\Omega)}\leq e^{-\mu t}\|\nabla e^{tD_{c}\Delta}c_{in}\|_{L^{\frac{n+1}{n+2}}(\Omega)}+Cr\int_{0}^{t}e^{-\mu(t-s)}(t-s)^{-\frac{n+1}{n+2}}\|d(s)\|_{L^{1}(\Omega)}\,ds.$
The first term is bounded uniformly on $[0,T]$ by the regularity of the initial datum, while the second term by (4.31). Therefore,
$\sup_{t\in[0,T]}\|\nabla c(t)\|_{L^{\frac{n+1}{n+2}}(\Omega)}\leq C,$
with $C$ independent of $\lambda$ and this concludes the proof.
We now extend the previous estimates to $L^{p}$ bounds for the solution components by means of a duality argument.
Lemma 12
Let $(h,d,m,c)\in\Phi$. Under the hypotheses of Theorem 5, for any $1\leq p<\infty$, there exists a constant $H_{p}$, depending on $p$, $\Omega$, $T$, $n$, the initial data and the parameters of the system, but not on $\lambda$, such that:
$\|h\|_{L^{p}(\Omega_{T})}+\|d\|_{L^{p}(\Omega_{T})}+\|m\|_{L^{p}(\Omega_{T})}\leq H_{p}\,.$ (4.36)
Proof. The proof follows [6, 13] (see also [3, 7, 8, 25, 28]).
We prove (4.36) for $p$ large enough, namely for $p=p{{}^{\prime}}/(p{{}^{\prime}}-1)$, where $p{{}^{\prime}}$ satisfies $1<p{{}^{\prime}}<\frac{n+2}{2}$. Consequently, we have $1+\frac{2}{n}<p<\infty$.
We proceed by a duality argument. Let $\theta\in L^{p{{}^{\prime}}}(\Omega_{T})$ be such that $\theta\geq 0$ and $\|\theta\|_{L^{p{{}^{\prime}}}(\Omega_{T})}\leq 1$. We denote by $\psi_{D}$ the unique solution of the following backward heat equation:
$\begin{cases}\begin{array}[]{l}\frac{\partial}{\partial t}\psi_{D}+D\Delta\psi_{D}=-\theta\qquad(t,x)\in\Omega_{T}\\ \nabla\psi_{D}\cdot\mathbf{n}=0\qquad\qquad(t,s)\in(0,T)\times\partial\Omega\\ \psi_{D}(T,x)=0\qquad\qquad x\in\Omega\end{array}\end{cases}$ (4.37)
The function $\psi_{D}$ is non-negative. Moreover, by maximal regularity for the heat semigroup [6, 22] and Lemma 6, we have:
$\|\partial_{t}\psi_{D}\|_{L^{p{{}^{\prime}}}(\Omega_{T})}+\|\psi_{D}\|_{L^{q_{1}}(\Omega_{T})}+\|\nabla\psi_{D}\|_{L^{q_{2}}(\Omega_{T})}\leq C_{p{{}^{\prime}}}\,$, (4.38)
where $q_{1}=(n+2)p{{}^{\prime}}/(n+2-2p{{}^{\prime}})$, $q_{2}=(n+2)p{{}^{\prime}}/(n+2-p{{}^{\prime}})$ and $C_{p{{}^{\prime}}}$ is a positive constant depending on $\Omega$, $T$, $n$ and $p{{}^{\prime}}$. Multiplying the equation by $\theta$, integrating over space and time and applying (4.37) with $D=D_{m}$ together with integration by parts and (4.30), yields:
$\int_{0}^{T}\int_{\Omega}m\,\theta\,dxdt\leq\int_{\Omega}m_{in}\psi_{D_{m}}(0,\cdot)\,dxdt+\int_{0}^{T}\int_{\Omega}\chi(m,c)\nabla c\nabla\psi_{D_{m}}\,dxdt\,\,+\\ +\nu\int_{0}^{T}\int_{\Omega}d\,\psi_{D_{m}}\,dxdt\,.$ (4.39)
Proceeding as above, multiplying $\theta$ by $h+d$, integrating over space and time, and applying (4.37) with $D=1$ and $D=D_{d}$ together with integration by parts and (4.30), yields:
$\int_{0}^{T}\int_{\Omega}(h+d)\,\theta\,dxdt\leq\int_{\Omega}h_{in}\,\psi_{1}(0,\cdot)\,dxdt+\int_{\Omega}d_{in}\,\psi_{D_{d}}(0,\cdot)\,dxdt+\rho\int_{0}^{T}\int_{\Omega}d\,\psi_{D_{d}}\,dxdt+\\ +\alpha(1+c_{\epsilon})\int_{0}^{T}\int_{\Omega}m\,h\,\psi_{D_{d}}\,dxdt\,\,+\int_{0}^{T}\int_{\Omega}(\sigma+\rho d)\,\psi_{1}\,dxdt\,.$ (4.40)
We now estimate each term in the above inequalities separately.
We begin with the initial term in (4.39). Using Hölder’s inequality together with (4.38) (see [6]), we obtain:
$\int_{\Omega}m_{in}\psi_{D_{m}}(0,x)dx\leq\|m_{in}\|_{L^{p}(\Omega)}\|\psi_{D_{m}}(0,x)\|_{L^{p{{}^{\prime}}}(\Omega)}\\ \leq\|m_{in}\|_{L^{p}(\Omega)}\,T^{1/p}\left\|\frac{\partial}{\partial_{t}}\psi_{D_{m}}\right\|_{L^{p{{}^{\prime}}}(\Omega_{T})}\leq\|m_{in}\|_{L^{p}(\Omega)}\,T^{1/p}\,C_{p{{}^{\prime}}}\,.$ (4.41)
The same argument applied to the first term in (4.40) gives:
$\int_{\Omega}h_{in}\psi_{1}(0,x)dx+\int_{\Omega}d_{in}\psi_{D_{d}}(0,x)dx\leq\left(\|h_{in}\|_{L^{p}(\Omega)}+\|d_{in}\|_{L^{p}(\Omega)}\right)\,T^{1/p}\,C_{p{{}^{\prime}}}\,.$ (4.42)
In the remainder of the proof, $C^{*}$ denotes a positive constant depending on $p$, $\Omega$, $T$ and $n$, but independent of $\lambda$.
We now estimate the last term in (4.40). Using (4.38), we have:
$\int_{0}^{T}\int_{\Omega}\sigma\psi_{1}\,dx\,dt=\sigma\|\psi_{1}\|_{L^{1}(\Omega_{T})}\leq\sigma C^{*}\|\psi_{1}\|_{L^{q_{1}}(\Omega_{T})}\leq\sigma C^{*}C_{p{{}^{\prime}}}\,.$ (4.43)
We next estimate the last term in (4.39). Following [6, 13], we introduce $\vartheta=1-\frac{2}{n+2}$, so that $0<\vartheta<1$. Using Hölder’s inequality together with (4.31), (4.38) and Lemma 6, we obtain:
$\int_{0}^{T}\int_{\Omega}d\psi_{D_{m}}\,dxdt\leq C^{\ast}\left(\int_{0}^{T}\int_{\Omega}\left(|d|^{1-\vartheta}\left|\psi_{D_{m}}\right|\right)^{p{{}^{\prime}}}\,dxdt\right)^{1/p{{}^{\prime}}}\,\|d\|^{\vartheta}_{L^{p}(\Omega_{T})}\\ \leq C^{\ast}\|\psi_{D_{m}}\|_{L^{q_{1}}(\Omega_{T})}\|d\|^{1-\vartheta}_{L^{1}(\Omega_{T})}\|d\|^{\vartheta}_{L^{p}(\Omega_{T})}\leq C^{*}C_{p{{}^{\prime}}}H_{1}^{1-\vartheta}\left\|d\right\|_{L^{p}(\Omega_{T})}^{\vartheta}\,.$ (4.44)
The same estimate applied to (4.40) yields:
$\int_{0}^{T}\int_{\Omega}d\,\left(\psi_{1}+\psi_{D_{d}}\right)\,dx\,dt\leq C^{*}C_{p{{}^{\prime}}}H_{1}^{1-\vartheta}\left\|d\right\|_{L^{p}(\Omega_{T})}^{\vartheta}\,.$ (4.45)
It remains to estimate the terms in (4.39) and (4.40).
We fix $s$ and $\xi$ in $(0,1)$ such that:
$(1-s)(n+2)<1\,,\qquad(2-\xi)(n+2)<1$ (4.46)
and consider:
$1<p{{}^{\prime}}<\min\left(\frac{n+2}{2}\,,\,\frac{2-(1-s)(n+2)}{n+2}\,,\,\frac{2-(2-\xi)(n+2)}{n+2}\right).$ (4.47)
Using (4.46) and (4.47), together with Hölder’s and Young’s inequalities and the estimates (4.31) and (4.35), we have:
$\int_{0}^{T}\int_{\Omega}\chi(m,c)\nabla c\nabla\psi_{D_{m}}\,dxdt\leq\chi_{0}\int_{0}^{T}\int_{\Omega}m\left|\nabla c\right|\left|\nabla\psi_{D_{m}}\right|\,dxdt\\ \leq\chi_{0}\left(\int_{0}^{T}\int_{\Omega}\left(|m|^{1-s}\left|\nabla c\right|\left|\nabla\psi_{D_{m}}\right|\right)^{p{{}^{\prime}}}\,dxdt\right)^{1/p{{}^{\prime}}}\,\|m\|^{s}_{L^{p}(\Omega_{T})}\\ \leq\chi_{0}\|\nabla\psi_{D_{m}}\|_{L^{q_{2}}(\Omega_{T})}\|m\|^{1-s}_{L^{1}(\Omega_{T})}\|\nabla c\|_{L^{\frac{n+2}{n+1}}(\Omega_{T})}\|m\|^{s}_{L^{p}(\Omega_{T})}\\ \leq\chi_{0}\,C_{p{{}^{\prime}}}H_{1}^{1-s}H^{\ast}\,\|m\|^{s}_{L^{p}(\Omega_{T})}\,$, (4.48)
and
[…]
Collecting the estimates (4.41), (4.44) and (4.48) in (4.39), together with (4.42), (4.45) and (4.49) in (4.40), we obtain:
$\int_{0}^{T}\int_{\Omega}(m+h+d)\,\theta\,dxdt\leq\\ \leq C_{1}+C_{2}\left\|m+h+d\right\|_{L^{p}(\Omega_{T})}^{\vartheta}+C_{3}\left\|m+h+d\right\|_{L^{p}(\Omega_{T})}^{s}+C_{4}\left\|m+h+d\right\|_{L^{p}(\Omega_{T})}^{\xi}\,$, (4.50)
where $C_{1}$, $C_{2}$, $C_{3}$ and $C_{4}$ are constants whose explicit expressions follow from the previous estimates and depend on $p$, $\Omega$, $T$, $n$, the initial data and the parameters of the system, but not on $\lambda$.
Since the above inequality holds for every $\theta\geq 0$ with $\|\theta\|_{L^{p{{}^{\prime}}}(\Omega_{T})}\leq 1$, duality in $L^{p}(\Omega_{T})$ yields:
$\left\|m+h+d\right\|_{L^{p}(\Omega_{T})}$ $\leq C_{1}+C_{2}\left\|m+h+d\right\|_{L^{p}(\Omega_{T})}^{\vartheta}+C_{3}\left\|m+h+d\right\|_{L^{p}(\Omega_{T})}^{s}$
$+C_{4}\left\|m+h+d\right\|_{L^{p}(\Omega_{T})}^{\xi}\,$,
with $\vartheta$, $s$ and $\xi\in(0,1)$.
The conclusion follows by an application of Young’s inequality.
We now show that the set $\Phi$ is bounded.
Lemma 13
Under the hypotheses of Theorem 5, the set $\Phi$ is bounded in $C^{0,1}]^{4}$. In particular, if $(h,d,m,c)\in\Phi$, then:
$\|(h,d,m,c)\|_{[C^{0,1}]^{4}}\leq C\,$, (4.51)
where $C$ is a constant depending on $\Omega$, $T$, $n$, the parameters of the system and the initial data, but not on $\lambda$.
Proof. Since $p>n+2$, the space $W^{1,2}_{p}(\Omega_{T})$ is compactly embedded in $C^{0,1}$ (see Lemma 6). Applying maximal regularity to the equation for $c$ in (4.30) and using (4.36), we obtain:
$\|c\|_{C^{0,1}}\leq C_{T}\|c\|_{W^{1,2}_{p}(\Omega_{T})}\leq C_{T}\left(r\|d\|_{L^{p}(\Omega_{T})}+\|c_{in}\|_{W^{p,2-2/p}(\Omega)}\right)\leq C\,$, (4.52)
where $C_{T}$ depends on $\Omega$, $T$, $n$, $p$, and $C$ depends on $\Omega$, $T$, $n$, $p$, the parameters of the system and the initial data, but not on $\lambda$.
Repeating the same argument for the $m$-equation in (4.30), we have:
$\|m\|_{C^{0,1}}\leq C_{T}\|m\|_{W^{1,2}_{p}(\Omega_{T})}\\ \leq C_{T}\left(\nu\|d\|_{L^{p}(\Omega_{T})}+\|\nabla\cdot\left(\chi(m,c)\nabla c\right)\|_{L^{p}(\Omega_{T})}+\|m_{in}\|_{W^{p,2-2/p}(\Omega)}\right)\leq C\,$, (4.53)
and analogously for the $d$-equation and the $h$-equation in (4.30):
$\|d\|_{C^{0,1}}\leq C_{T}\|d\|_{W^{1,2}_{p}(\Omega_{T})}\\ \leq C_{T}\left(\|A(m,c)h\|_{L^{p}(\Omega_{T})}+\rho\lambda\|d\overline{(h+d)}_{+}\|_{L^{p}(\Omega_{T})}+\|d_{in}\|_{W^{p,2-2/p}(\Omega)}\right)\leq C\,$, (4.54)
$\|h\|_{C^{0,1}}\leq C_{T}\|h\|_{W^{1,2}_{p}(\Omega_{T})}\\ \leq C_{T}\left(\|(\lambda\sigma+\rho d)(1-\overline{(h+d)}_{+})\|_{L^{p}(\Omega_{T})}+\|h_{in}\|_{W^{p,2-2/p}(\Omega)}\right)\leq C\,.$ (4.55)
Combining (4.55), (4.54), (4.53) and (4.52), we obtain (4.51).
We conclude by proving that the saturation condition for $h$ and $d$ is preserved.
Lemma 14
Let $(h,d,m,c)\in\Phi$. Then:
$h+d\leq 1.$
Proof. From (4.30), the $2\times 2$ parabolic semilinear system for $h$ and $d$ has a diagonal diffusion. Arguing as in Theorem 1 and considering Lemma 9, the reaction field is inward pointing on the boundary face $h+d=1$, and on $h=0$ and $d=0$. Applying the invariant-region theorem for semilinear parabolic systems (see [4, 34]) yields that the convex bounded domain
$\Sigma=\{(h,d)\in\mathbb{R}^{2}:\ h\geq 0,\ d\geq 0,\ h+d\leq 1\}$
is positively invariant and the Lemma is proved.
By Lemma 14, we have $h+d\leq 1$. In particular, the truncation $\overline{(h+d)}_{+}$ coincides with $(h+d)$ and system (4.27) reduces to:
$\begin{cases}\displaystyle\partial_{t}h=\Delta h-h-A(m,c)h+\lambda\sigma(1-(h+d))+\rho d(1-(h+d)),&(t,x)\in\Omega_{T}\\ \displaystyle\partial_{t}d=D_{d}\Delta d-(\delta+\rho)d+A(m,c)h+\rho\lambda d(h+d),&(t,x)\in\Omega_{T}\\ \displaystyle\partial_{t}m=D_{m}\Delta m-\nabla\cdot\left(\chi(m,c)\nabla c\right)+\nu(\lambda d-m),&(t,x)\in\Omega_{T}\\ \displaystyle\partial_{t}c=D_{c}\Delta c-\mu c+r\lambda d\,,&(t,x)\in\Omega_{T}\end{cases}$ (4.56)
with boundary conditions (4.28) and nonnegative initial data (4.29) with condition (4.5).
4.1.5 Existence of the solution
In this subsection, we prove the existence of a global solution. The result is stated in the following proposition, whose proof follows from Lemmas 9, 7, 8, 13 and 14.
Proposition 15
Under the hypotheses of Theorem 5, for any $T>0$, the system (4.56) with Neumann boundary conditions (4.3) and non-negative initial data (4.4) satisfying (4.5), admits a strong non-negative solution $(h,d,m,c)$ in $[0,T]$.
Moreover, there exists a constant $C$, depending on $\Omega$, $T$, $n$, the parameters of the system and $\|(h_{in},d_{in},m_{in},c_{in})\|_{[W^{2-2/\bar{p},\bar{p}}(\Omega)]^{4}}$, such that:
$\|(h,d,m,c)\|_{[C^{0,1}]^{4}}\leq C\,.$ (4.57)
Proof. By the Leray–Schauder fixed point theorem, the map $S$ admits a fixed point.
This yields, for any $T>0$, a solution of the original system (4.1) in $[0,T]$.
Estimate (4.57) follows from (4.51).
Finally, since all terms in (4.1) are defined almost everywhere as $L^{1}(\Omega_{T})$ functions, the solution is strong.
4.2 Uniqueness
In this subsection, we prove the uniqueness of the solution to system (4.1) with boundary conditions (4.3) and non-negative initial data (4.4) satisfying (4.5).
Proposition 16
Under the hypotheses of Theorem 5, for each $T>0$, the solution to system (4.1) with Neumann boundary conditions (4.3) and non-negative initial data (4.4) satisfying (4.5), constructed in Proposition 15, is unique.
Proof. Let $(h_{1},d_{1},m_{1},c_{1})$ and $(h_{2},d_{2},m_{2},c_{2})$ be two solutions of system (4.1), with boundary and initial conditions given in (4.3) and (4.4) with (4.5). Taking their differences and using Young inequality, we have:
$\frac{d}{dt}\|c_{1}-c_{2}\|^{2}_{L^{2}(\Omega)}+2D_{c}\|\nabla(c_{1}-c_{2})\|^{2}_{L^{2}(\Omega)}\leq r\|d_{1}-d_{2}\|^{2}_{L^{2}(\Omega)}+r\|c_{1}-c_{2}\|^{2}_{L^{2}(\Omega)}-2\mu\|c_{1}-c_{2}\|^{2}_{L^{2}(\Omega)}\,$, (4.58)
$\frac{d}{dt}\|m_{1}-m_{2}\|^{2}_{L^{2}(\Omega)}+2D_{m}\|\nabla(m_{1}-m_{2})\|^{2}_{L^{2}(\Omega)}\leq\nu\|d_{1}-d_{2}\|^{2}_{L^{2}(\Omega)}-\nu\|m_{1}-m_{2}\|^{2}_{L^{2}(\Omega)}+\\ \quad\quad\quad+\frac{3\chi_{0}^{2}}{2D_{m}}\|m_{1}\|^{2}_{C^{0,1}}\|\nabla(c_{1}-c_{2})\|^{2}_{L^{2}(\Omega)}+\frac{2}{3}D_{m}\|\nabla(m_{1}-m_{2})\|^{2}_{L^{2}(\Omega)}+\\ \quad\quad\quad+\frac{3\chi_{0}^{2}}{2D_{m}}\|c_{2}\|^{2}_{C^{0,1}}\|m_{1}-m_{2})\|^{2}_{L^{2}(\Omega)}+\frac{2}{3}D_{m}\|\nabla(m_{1}-m_{2})\|^{2}_{L^{2}(\Omega)}+\\ \quad\quad\quad+\frac{3\chi_{0}^{2}}{2\kappa^{2}D_{m}}\|m_{2}\|^{2}_{C^{0,1}}\|c_{2}\|^{2}_{C^{0,1}}\|c_{1}-c_{2}\|^{2}_{L^{2}(\Omega)}+\frac{2}{3}D_{m}\|\nabla(m_{1}-m_{2})\|^{2}_{L^{2}(\Omega)}\,$, (4.59)
[…]
Denoting by $A=\frac{3}{4}\frac{\chi_{0}^{2}}{D_{m}\,D_{c}}\|m_{1}\|^{2}_{C^{0,1}}$, we complete the proof applying the Gronwall’s lemma to:
$\|h_{1}-h_{2}\|^{2}_{L^{2}(\Omega)}+\|d_{1}-d_{2}\|^{2}_{L^{2}(\Omega)}+\|m_{1}-m_{2}\|^{2}_{L^{2}(\Omega)}+A\|c_{1}-c_{2}\|^{2}_{L^{2}(\Omega)}\,$,
using the previous estimates (4.58)-(4.59)-(4.60) and estimate (4.57) given in Proposition 15.
4.3 Uniform-in-time estimates
In this section we will conclude the proof of Theorem 5 proving (4.6) and (4.7).
We prove three preliminary Lemmas where we establish, respectively, uniform in time bounds in $L^{1}(\Omega)$ for the solution, uniform in time bounds for the gradient of $c$ and finally we state uniform in time bounds in $L^{p}(\Omega)$ for the solution. We conclude this section with Proposition 20, where we prove the uniform in time estimate of the solution in $W^{1,\infty}(\Omega)$ (see [6]).
Lemma 17
Under the hypothesis of Theorem 5, let $(h,d,m,c)$ be the unique solution to (4.1) on $\mathbb{R}_{+}\times\Omega$, then there exists a constant $\overline{H}_{1}$, depending on $\Omega$, $n$, the parameters of the system and the initial data, such that:
$\sup_{t\in\mathbb{R}_{+}}\left(\|(h,d,m,c)\|_{[L^{1}(\Omega)]^{4}}\right)\leq\overline{H}_{1}\,$, (4.61)
and
$\limsup_{t\rightarrow\infty}\left(\|(h,d,m,c)\|_{[L^{1}(\Omega)]^{4}}\right)\leq\overline{H}_{1}\,.$ (4.62)
Proof. Integrating the equations for $h$ and $d$ in (4.1), summing and considering the saturation condition given in Lemma 14, we obtain:
$\partial_{t}\int_{\Omega}(h+d)dx\leq\sigma\int_{\Omega}(1-(h+d))dx\,$, (4.63)
Applying Gronwall’s lemma, we have (4.61) and (4.62) for $h$ and $d$.
Integrating the equation of $m$ and $c$ in (4.1), we obtain:
$\frac{d}{dt}\int_{\Omega}m\,dx\leq-\nu\int_{\Omega}m\,dx+\nu\|d\|_{L^{1}(\Omega)}\,$,
and
$\frac{d}{dt}\int_{\Omega}c\,dx\leq-\mu\int_{\Omega}c\,dx+r\|d\|_{L^{1}(\Omega)}\,.$
Moreover, (4.61) and (4.62) for $m$ and $c$ are proved applying Gronwall’s lemma.
To prove the next two lemmas, following [6], we define a smooth positive cutoff function $\upsilon\in C^{\infty}\left(\mathbb{R};[0,1]\right)$ such that $\sup_{s\in\mathbb{R}}|\upsilon(s)|\leq 1$ and $\upsilon(s)=1$ for $s\geq 1$ and $\upsilon(s)=0$ for $s\leq 0$ and $|\upsilon{{}^{\prime}}(s)|\leq 2$ for all $s\in\mathbb{R}$. For any $\tau\in\mathbb{N}$, we define $\upsilon_{\tau}(\cdot)=\upsilon(\cdot-\tau)$. The idea is to multiply the equations (4.1) by $\upsilon_{\tau}$ and repeating the steps used in Lemmas 11 and 12 to obtain new estimates in $\Omega_{\tau,\tau+1}$ which are independent of $\tau$.
Lemma 18
Under the hypothesis of Theorem 5, let $(h,d,m,c)$ be the unique solution to (4.1) on $\mathbb{R}_{+}\times\Omega$, then there exists a constant $\overline{H}^{\ast}$, depending on $\Omega$, $n$, the parameters of the system and the initial data, but not on $\tau$, such that:
$\sup_{\tau\in\mathbb{N}-\{0\}}\|\nabla c\|_{L^{\frac{n+2}{n+1}}(\Omega_{\tau,\tau+1})}\leq\overline{H}^{\ast}\,$, (4.64)
and
$\limsup_{\tau\rightarrow\infty}\|\nabla c\|_{L^{\frac{n+2}{n+1}}(\Omega_{\tau,\tau+1})}\leq\overline{H}^{\ast}\,.$ (4.65)
Proof.
Repeating the same arguments in Lemma 11 for the equation of $\upsilon_{\tau}\,c$, considering that $\upsilon_{\tau}c=0$ in $(x,\tau)$ and using (4.61) and (4.62) and the properties of $\upsilon_{\tau}$, we obtain (4.64) and (4.65).
Lemma 19
Under the hypothesis of Theorem 5, let $(h,d,m,c)$ be the unique solution to (4.1) on $\mathbb{R}_{+}\times\Omega$, then there exists a constant $\overline{H}_{p}$, depending on $\Omega$, $p$, $n$, the parameters of the system and the initial data, but not on $\tau$, such that:
$\sup_{\tau\in\mathbb{N}-\{0\}}\left(\|(h,d,m,c)\|_{[L^{p}(\Omega_{\tau,\tau+1})]^{4}}\right)\leq\overline{H}_{p}\,$, (4.66)
and
$\limsup_{\tau\rightarrow\infty}\left(\|(h,d,m,c)\|_{[L^{p}(\Omega_{\tau,\tau+1})]^{4}}\right)\leq\overline{H}_{p}\,.$ (4.67)
for any $1\leq p<\infty$.
Proof. The proof follows the duality arguments performed in Lemma 12. The main difference is that the equations for $h$, $d$, $m$ and $c$ are multiplied by $\upsilon_{\tau}$ and equation (4.37) and regularity estimates (4.38) are considered in the domain $\Omega_{\tau,\tau+2}$.
Multiplying $\theta$ by $\upsilon_{\tau}m$, integrating in space and time and using equation (4.37) in $\Omega_{\tau,\tau+2}$, integrating thus by parts and using (4.1), we obtain:
$\int_{\tau}^{\tau+2}\int_{\Omega}\upsilon_{\tau}m\,\theta\,dxdt\leq\int_{\tau}^{\tau+2}\int_{\Omega}\chi(m,c)\nabla c\nabla\psi_{D_{m}}\,dxdt\,\,+\\ +\int_{\tau}^{\tau+2}\int_{\Omega}\upsilon{{}^{\prime}}_{\tau}m\,\psi_{D_{m}}\,dxdt+\nu\int_{\tau}^{\tau+2}\int_{\Omega}\upsilon_{\tau}d\,\psi_{D_{m}}\,dxdt\,.$ (4.68)
Proceeding as above, we multiply $\theta$ by $\upsilon_{\tau}(h+d)$, integrate over space and time, and apply (4.37) in $\Omega_{\tau,\tau+2}$ together with integration by parts and (4.1), which yields:
$\int_{\tau}^{\tau+2}\int_{\Omega}\upsilon_{\tau}(h+d)\,\theta\,dxdt\leq\alpha(1+c_{\epsilon})\int_{\tau}^{\tau+2}\int_{\Omega}m\,h\,\psi_{D_{d}}\,dxdt\,\,+\rho\int_{\tau}^{\tau+2}\int_{\Omega}\upsilon_{\tau}d\,\psi_{D_{d}}\,dxdt+\\ +\int_{\tau}^{\tau+2}\int_{\Omega}\upsilon_{\tau}(\sigma+\rho d)\,\psi_{1}\,dxdt+\int_{\tau}^{\tau+2}\int_{\Omega}\upsilon{{}^{\prime}}_{\tau}d\psi_{D_{d}}\,dxdt+\int_{\tau}^{\tau+2}\int_{\Omega}\upsilon{{}^{\prime}}_{\tau}h\psi_{1}\,dxdt\,.$ (4.69)
For the equation of $c$, we have:
$\int_{\tau}^{\tau+2}\int_{\Omega}\upsilon_{\tau}\,c\,\theta\,dxdt\leq r\int_{\tau}^{\tau+2}\int_{\Omega}\upsilon_{\tau}d\,\psi_{D_{c}}\,dxdt+\int_{\tau}^{\tau+2}\int_{\Omega}\upsilon{{}^{\prime}}_{\tau}c\,\psi_{D_{c}}\,dxdt\,.$ (4.70)
We can estimate each terms in (4.68), (4.69) and (4.70) following the proof in Lemma 12 and using (4.61), (4.62), (4.64) and (4.65), obtaining:
$\left\|\upsilon_{\tau}(m+h+d+c)\right\|_{L^{p}(\Omega_{\tau,\tau+2})}\leq\overline{C}_{1}+\overline{C}_{2}\left\|\upsilon_{\tau}(m+h+d+c)\right\|_{L^{p}(\Omega_{\tau,\tau+2})}^{\vartheta}\\ +\overline{C}_{3}\left\|\upsilon_{\tau}(m+h+d+c)\right\|_{L^{p}(\Omega_{\tau,\tau+2})}^{s}+\overline{C}_{4}\left\|\upsilon_{\tau}(m+h+d+c)\right\|_{L^{p}(\Omega_{\tau,\tau+2})}^{\xi}\,$,
with $\vartheta$, $s$ and $\xi\in(0,1)$. The constants $\overline{C}_{1}$, $\overline{C}_{2}$, $\overline{C}_{3}$ and $\overline{C}_{4}$ can be computed explicitly and depend on $\Omega$, $n$, $p$, the initial data and the parameters of the system, but not on $\tau$. We obtain (4.66) and (4.67) using Young’s inequality and the properties of $\upsilon_{\tau}$.
We conclude with the following proposition.
Proposition 20
Under the hypotheses of Theorem 5, let $(h,d,m,c)$ be the unique solution to system (4.1) with Neumann boundary conditions (4.3) and initial data (4.4), constructed in Propositions 15 and 16. Then, there exists a constant $\overline{C}>0$, depending on $\Omega$, $h_{in}$, $d_{in}$, $m_{in}$, $c_{in}$ and on the parameters of the system, such that the solution satisfies:
$\sup_{t\geq 0}\left\|\left(h(t,\cdot)\,,\,d(t,\cdot)\,,\,m(t,\cdot)\,,\,c(t,\cdot)\right)\right\|_{[L^{\infty}(\Omega)]^{4}}\leq C\,$, (4.71)
$\limsup_{t\rightarrow+\infty}\left\|\left(h(t,\cdot)\,,\,d(t,\cdot)\,,\,m(t,\cdot)\,,\,c(t,\cdot)\right)\right\|_{[L^{\infty}(\Omega)]^{4}}\leq C\,$, (4.72)
and
$\sup_{t\geq 0}\left\|\left(\nabla h(t,\cdot)\,,\,\nabla d(t,\cdot)\,,\,\nabla m(t,\cdot)\,,\,\nabla c(t,\cdot)\right)\right\|_{[L^{\infty}(\Omega)]^{4}}\leq C\,$, (4.73)
$\limsup_{t\rightarrow+\infty}\left\|\left(\nabla h(t,\cdot)\,,\,\nabla d(t,\cdot)\,,\,\nabla m(t,\cdot)\,,\,\nabla c(t,\cdot)\right)\right\|_{[L^{\infty}(\Omega)]^{4}}\leq C\,.$ (4.74)
Proof. Using the embedding properties stated in Lemma 6 together with the maximal regularity theory for the heat equation and taking $p>n+2$, we obtain:
$\|\upsilon_{\tau}c\|_{L^{\infty}(\Omega_{\tau,\tau+2})}+\|\upsilon_{\tau}\nabla c\|_{L^{\infty}(\Omega_{\tau,\tau+2})}\leq\overline{C}^{\ast}\|\upsilon_{\tau}c\|_{W_{p}^{1,2}(\Omega_{\tau,\tau+2})}$
$\leq\overline{C}^{\ast}\|\upsilon_{\tau}{{}^{\prime}}c-\upsilon_{\tau}\mu c+r\upsilon_{\tau}\,d\|_{L^{p}(\Omega_{\tau,\tau+2})}\,$, (4.75)
where $\overline{C}^{\ast}$ depends on $\Omega$, $n$ and the parameters of the system, but not on $\tau$. Using (4.66), (4.67) and the definition of $\upsilon_{\tau}$, the Lemma is proved for $c$.
Analogously for $h$ and $d$, as:
$\|\upsilon_{\tau}h\|_{L^{\infty}(\Omega_{\tau,\tau+2})}+\|\upsilon_{\tau}\nabla h\|_{L^{\infty}(\Omega_{\tau,\tau+2})}\leq\overline{C}^{\ast}\|\upsilon_{\tau}h\|_{W_{p}^{1,2}(\Omega_{\tau,\tau+2})}\leq$
$\overline{C}^{\ast}\|(\upsilon_{\tau}{{}^{\prime}}-\upsilon_{\tau})h\|_{L^{p}(\Omega_{\tau,\tau+2})}+\overline{C}^{\ast}\|\upsilon_{\tau}(\sigma+\rho\,d)(1-h-d)-\upsilon_{\tau}\,A(m,c)h\|_{L^{p}(\Omega_{\tau,\tau+2})}\,$
and
$\|\upsilon_{\tau}d\|_{L^{\infty}(\Omega_{\tau,\tau+2})}+\|\upsilon_{\tau}\nabla d\|_{L^{\infty}(\Omega_{\tau,\tau+2})}\leq\overline{C}^{\ast}\|\upsilon_{\tau}d\|_{W_{p}^{1,2}(\Omega_{\tau,\tau+2})}\leq$
$\overline{C}^{\ast}\|(\upsilon_{\tau}{{}^{\prime}}-\delta\upsilon_{\tau})d\|_{L^{p}(\Omega_{\tau,\tau+2})}+\overline{C}^{\ast}\|\upsilon_{\tau}\,A(m,c)h-\upsilon_{\tau}\,\rho\,d(1-h-d)\|_{L^{p}(\Omega_{\tau,\tau+2})}\,.$
For the $m$-equation, following [6], we write:
$\upsilon_{\tau}m=\int_{\tau}^{t}e^{-D_{m}\Delta(t-\tau-s)}\left(\upsilon_{\tau}\nu(d-m)+\upsilon_{\tau}{{}^{\prime}}m-\upsilon_{\tau}\nabla\cdot\left(\chi(m,c)\nabla c\right)\right)ds\,$, (4.76)
and
$\upsilon_{\tau}\nabla m=\int_{\tau}^{t}e^{-D_{m}\Delta(t-\tau-s)}\nabla\left(\upsilon_{\tau}\nu(d-m)+\upsilon_{\tau}{{}^{\prime}}m-\upsilon_{\tau}\nabla\cdot\left(\chi(m,c)\nabla c\right)\right)ds\,.$ (4.77)
Using the $L^{p}-L^{q}$ estimate (see [14]) of the heat kernel, we have:
$\sup_{\tau<t<\tau+2}$ $\|\upsilon_{\tau}m\|_{L^{\infty}(\Omega)}\leq$
$\overline{C}^{\ast}\left(\|m\|_{L^{p}(\Omega_{\tau,\tau+2})}+\|d\|_{L^{p}(\Omega_{\tau,\tau+2})}\right)\int_{\tau}^{\tau+2}\left(1+(t-\tau-s)^{-n/2p}\right)ds+$
$+\overline{C}^{\ast}\chi_{0}\|\nabla\,c\|_{L^{\infty}(\Omega_{\tau,\tau+2})}\|m\|_{L^{p}(\Omega_{\tau,\tau+2})}\int_{\tau}^{\tau+2}\left(1+(t-\tau-s)^{-\frac{1}{2}-\frac{n}{2p}}\right)ds\,$, (4.78)
and
$\sup_{\tau<t<\tau+2}$ $\|\upsilon_{\tau}\nabla m\|_{L^{\infty}(\Omega)}\leq$
$\overline{C}^{\ast}\left(\|m\|_{L^{p}(\Omega_{\tau,\tau+2})}+\|d\|_{L^{p}(\Omega_{\tau,\tau+2})}\right)\int_{\tau}^{\tau+2}\left(1+(t-\tau-s)^{-\frac{1}{2}-\frac{n}{2p}}\right)ds+$
$+\overline{C}^{\ast}\chi_{0}\|\nabla\,c\|_{L^{\infty}(\Omega_{\tau,\tau+2})}\|m\|_{L^{p}(\Omega_{\tau,\tau+2})}\int_{\tau}^{\tau+2}\left(1+(t-\tau-s)^{-1-\frac{n}{2p}}\right)ds\,$, (4.79)
where $\overline{C}^{\ast}$ depends on $\Omega$, $n$ and the parameters of the system, but not on $\tau$.
Choosing $p>n$ and using (4.66), (4.67) together with the properties of $\upsilon_{\tau}$, we obtain (4.73) and (4.74) for $m$.   Proposition 15, Proposition 16 and Proposition 20 complete the proof of Theorem 5.
The results of this section provide the analytical framework ensuring that the model is well-posed and biologically consistent, which allows for the investigation of the dynamical properties of the system in the next section.
5 Early-stage dynamics: invasion and propagation
In this section, we investigate the onset of disease spreading, focusing on the early stage of invasion where damage and inflammatory activity remain small. In this regime, the dynamics is governed by small perturbations of the healthy equilibrium and can therefore be analyzed by linearization.
Our goal is to identify the mechanisms responsible for the initiation of spatial propagation. In particular, we determine whether spatial heterogeneity may arise through diffusion-driven instability, or whether disease progression occurs through invasion processes.
To this end, we linearize the system around the healthy equilibrium and derive the associated dispersion relation. This allows us to characterize both the invasion threshold and the propagation speed of pathological fronts in the early-stage regime.
In order to linearize the system around the healthy equilibrium $H^{*}$ given in (3.3), we introduce the small perturbations:
$\tilde{h}=h_{0}+h,\qquad\tilde{d}=d,\qquad\tilde{m}=m,\qquad\tilde{c}=c.$ (5.1)
Substituting in (2.3), the resulting linearized system is:
$\dot{\mathbf{w}}=K\mathbf{w}+D\Delta\mathbf{w},\qquad\mathrm{with}\quad\mathbf{w}=\begin{pmatrix}\tilde{h}-h_{0}\\ \tilde{d}\\ \tilde{m}\\ \tilde{c}\end{pmatrix}$ (5.2)
and
$K=\begin{pmatrix}-(1+\sigma)&\dfrac{\rho}{1+\sigma}-\sigma&-\alpha c_{\epsilon}h^{\star}&0\\ 0&-\delta-\dfrac{\rho}{1+\sigma}&\alpha c_{\epsilon}h^{\star}&0\\ 0&\nu&-\nu&0\\ 0&r&0&-\mu\end{pmatrix},\qquad D=\begin{pmatrix}1&0&0&0\\ 0&D_{d}&0&0\\ 0&0&D_{m}&0\\ 0&0&0&D_{c}\end{pmatrix}$. (5.3)
5.1 Absence of diffusion-driven instabilities
We first investigate whether spatial heterogeneity can arise through Turing instability. Such mechanisms would lead to pattern formation from small perturbations of the homogeneous equilibrium.
To this end, we consider normal mode solutions of the form $e^{ikx+\lambda t}$. Substituting this ansatz into (5.2), the eigenvalues $\lambda$ are determined by the characteristic equation:
$|\lambda I-\Gamma K+k^{2}D|=0,$
i.e.
$(\lambda+D_{c}k^{2}+\mu)(\lambda+\sigma+1+k^{2})\left[\lambda^{2}+\left(\delta+D_{d}k^{2}+D_{m}k^{2}+\nu+\dfrac{\rho}{1+\sigma}\right)\lambda+h(k^{2})\right]=0,$ (5.4)
where
$h(k^{2})=D_{d}D_{m}k^{4}+\left(\delta D_{m}+D_{d}\nu+D_{m}\dfrac{\rho}{1+\sigma}\right)k^{2}+\nu\delta+\nu\dfrac{\rho}{1+\sigma}-\nu\alpha c_{\epsilon}\dfrac{\sigma}{1+\sigma}.$ (5.5)
Spatially heterogeneous modes correspond to wavenumbers $k$ such that $\mathrm{Re}(\lambda)>0$. Therefore, diffusion-driven instability can only occur if there exists $k>0$ such that $h(k^{2})<0$. Marginal stability is reached when the minimum of $h(k^{2})$ vanishes. This condition yields the critical value:
$k^{2}=-\dfrac{1}{D_{d}D_{m}}\left(D_{m}\delta+D_{m}\dfrac{\rho}{1+\sigma}+D_{d}\nu\right).$ (5.6)
Since the right-hand side in (5.6) is negative for all admissible parameter values, no real wavenumber $k$ satisfies this condition. Consequently, diffusion-driven (Turing) instabilities cannot occur in this system.
This result shows that spatial pattern formation cannot be triggered by diffusion alone. In particular, the model does not support the emergence of patchy or heterogeneous spatial structures arising from Turing-type mechanisms. Therefore, disease progression cannot occur through the spontaneous formation of localized damaged regions within otherwise healthy tissue. Instead, any spatial spreading of damage must arise through invasion processes, in which localized perturbations propagate across the tissue in the form of traveling fronts, as analyzed in the following subsection.
5.2 Invasion threshold and propagation speed
In this subsection, we characterize the conditions under which localized damage is able to invade healthy tissue and propagate across the domain. Based on the linearization around the healthy equilibrium, we derive both the invasion threshold and the propagation speed of the resulting fronts.
This analysis complements the previous result on the absence of diffusion-driven instabilities: since pattern formation cannot arise through Turing mechanisms, spatial progression must be driven by invasion processes. The linearized system allows us to explicitly identify the parameter regime in which small perturbations grow and initiate spatial spreading.
As a first step, we examine the temporal stability of the system in the long-wavelength limit ($k=0$). In this regime, the characteristic polynomial factorizes as follows:
$p(\lambda)=(\lambda+\mu)(1+\sigma+\lambda)\tilde{g}(\lambda)=0,$ (5.7)
and the dynamics are governed by the quadratic factor:
$\tilde{g}(\lambda)=\lambda^{2}+\left(\delta+\nu+\dfrac{\rho}{1+\sigma}\right)\lambda+\nu\delta+\nu\dfrac{\rho}{1+\sigma}-\nu\alpha c_{\epsilon}\dfrac{\sigma}{1+\sigma}.$ (5.8)
Our objective is to determine whether $\tilde{g}(\lambda)=0$ admits a positive real root, which indicates instability of the healthy state and, consequently, the onset of invasion.
This equation admits two real roots, as shown by the discriminant
$\Delta=\left(\delta-\nu+\dfrac{\rho}{1+\sigma}\right)^{2}+4\dfrac{\alpha c_{\epsilon}\nu\sigma}{1+\sigma}>0,$ (5.9)
which is strictly positive. By Descartes’ rule of signs, the equation admits at least one negative root. A unique positive root exists if and only if the following threshold condition holds:
$\delta(1+\sigma)+\rho-\alpha c_{\epsilon}\sigma<0.$ (5.10)
This condition identifies the parameter regime in which the homogeneous mode becomes linearly unstable. In particular, when (5.10) holds, the quadratic factor $\tilde{g}(\lambda)$ admits a positive root, so that small perturbations of the healthy state grow exponentially in time, leading to the onset of invasion in the early-stage regime.
Remark 21 (Interpretation of the invasion threshold.)
The invasion condition can be interpreted in terms of an effective damage reproduction number. Rewriting the threshold condition:
$\delta(1+\sigma)+\rho-\alpha c_{\epsilon}\sigma=0$
yields:
$R_{d}=\frac{\alpha c_{\epsilon}\sigma}{\delta(1+\sigma)+\rho}.$ (5.11)
This quantity measures the balance between damage amplification driven by immune activity (the term $\alpha c_{\epsilon}\sigma$) and the mechanisms responsible for damage removal and tissue repair (the term $\delta(1+\sigma)+\rho$). Spatial invasion becomes possible when $R_{d}>1$, whereas for $R_{d}<1$ small perturbations of the healthy state decay. Biologically, this threshold expresses the transition between a regime in which repair mechanisms dominate and suppress damage and a regime in which immune-mediated damage overcomes tissue recovery, allowing degeneration to spread.
To characterize this invasion, we look for traveling front solutions of the linearized system in the form:
$\mathbf{w}(x,t)=\mathbf{v}e^{-\gamma(x-st)},$ (5.12)
where $s>0$ represents the propagation speed, $\gamma>0$ is the spatial decay rate of the front and $\mathbf{v}=(\hat{h},\hat{d},\hat{m},\hat{c})^{T}$ is the amplitude vector. Substituting this ansatz into the linearized system, we obtain the following eigenvalue problem:
$(\gamma^{2}D+K-\gamma s\mathbb{I})\mathbf{v}=0.$ (5.13)
Non-trivial solutions exist only when $\det(\gamma^{2}D+K-\gamma s\mathbb{I})=0$. Due to the block structure of the system matrices, the determinant factorizes as:
$(\gamma^{2}D_{c}-\gamma s-\mu)(\gamma^{2}-\gamma s-(1+\sigma))p_{2}(\gamma,s)=0,$ (5.14)
where:
$p_{2}(\gamma,s)=\gamma^{2}s^{2}+\gamma\left[\delta-(D_{d}+D_{m})\gamma^{2}+\nu+\dfrac{\rho}{1+\sigma}\right]s+D_{d}D_{m}\gamma^{4}$ (5.15)
$+\gamma^{2}\left(-D_{m}\dfrac{\rho}{1+\sigma}-D_{m}\delta-\nu D_{d}\right)+\underbrace{\delta\nu+\nu\dfrac{\rho}{1+\sigma}-\dfrac{\alpha c_{\epsilon}\nu\sigma}{1+\sigma}}_{\mathcal{C}}.$
The first two factors in (5.14) correspond to modes associated with the decoupled dynamics of the chemokine and healthy tissue components. These modes do not generate instability and therefore do not contribute to invasion.
In contrast, the quadratic factor $p_{2}(\gamma,s)$ captures the coupled dynamics of damaged tissue and immune response, which are responsible for the onset of instability and spatial spreading. For this reason, the propagation speed of the invasion front is determined by the roots of $p_{2}(\gamma,s)=0$. For fixed $\gamma>0$, this is a quadratic equation in $s$, whose discriminant is:
$\Delta(p_{2})=\gamma^{2}\left(\delta-D_{d}\gamma^{2}+D_{m}\gamma^{2}-\nu+\dfrac{\rho}{1+\sigma}\right)^{2}+4\alpha c_{\epsilon}\gamma^{2}\nu\dfrac{\sigma}{1+\sigma}>0.$ (5.16)
Hence, for any $\gamma>0$, the equation admits two real roots.
A physically meaningful invasion requires the existence of a positive propagation speed ($s>0$). To establish this, we analyze the constant term $\mathcal{C}$ of $p_{2}$, which is negative under the invasion condition (5.10). In this case, the polynomial $p_{2}(\gamma,0)$, viewed as a quadratic function of $z=\gamma^{2}$, admits a unique positive root $z^{\star}$. Consequently, there exists a corresponding value $\gamma^{\star}>0$ such that:
$p_{2}(\gamma,0)<0\qquad\text{for}\quad 0<\gamma<\gamma^{\star}.$
For such values of $\gamma$, the equation $p_{2}(\gamma,s)=0$, regarded as a quadratic polynomial in $s$ (for fixed $\gamma$), has a negative constant term and positive leading coefficient and therefore admits exactly one positive root. This defines a positive branch $s_{+}(\gamma)$ corresponding to admissible propagation speeds.
The biologically relevant front speed is then obtained as the minimal value of this branch:
$s^{*}=\min_{\gamma>0}s_{+}(\gamma),$ (5.17)
whenever the minimum is attained. This selection criterion corresponds to the classical linear spreading speed, which characterizes invasion fronts governed by the growth of small perturbations at the leading edge. In analogy with Fisher-KPP type systems, this suggests that the propagation is driven by a pulled front mechanism [39]. This interpretation will be corroborated by numerical simulations of the full nonlinear model in section 6.
Remark 22 (Biological interpretation of the propagation speed)
The minimal speed $s^{*}$ quantifies the rate at which damage spreads across the tissue once invasion is triggered. In biological terms, it represents the velocity at which degenerative regions expand through the muscle.
This highlights that, even when the healthy state becomes unstable, disease progression is not instantaneous but occurs through a spatially organized process, with a well-defined propagation speed determined by the balance between damage amplification and tissue repair processes.
The above analysis provides a coherent description of the early spatial dynamics of DMD. The invasion threshold identifies the parameter regime in which localized damage can grow from small perturbations of the healthy state, while the associated propagation speed quantifies the rate at which degenerative regions expand through the tissue.
In this early-stage regime, propagation is determined by the linearized dynamics near the front, leading to a propagation mechanism consistent with a pulled front.
Remark 23 (Role of chemotaxis)
The chemotactic flux does not affect the invasion threshold or the linear spreading speed. Indeed, the chemotaxis term involves nonlinear contributions of the form $m\nabla c$ and therefore vanishes in the linearization around the healthy equilibrium.
This shows that, in the early-stage regime considered in this work, the onset of spatial disease propagation is entirely controlled by the local damage–inflammation feedback, independently of directed immune migration. In particular, invasion is driven by reaction mechanisms rather than by chemotactic effects.
As a consequence, chemotaxis has only a limited quantitative influence on the early invasion dynamics and does not significantly alter the shape or speed of the propagating fronts, as confirmed by numerical simulations. Its inclusion nevertheless provides a structurally consistent framework for the investigation of later stages of the disease, where stronger inflammatory activity may lead to non-negligible chemotactic effects and more complex spatial behaviors.
6 Numerical simulations
In this section, we investigate the dynamical behavior of the model through numerical simulations of the full nonlinear system. The aim is twofold: first, to validate the analytical predictions derived from the linearized system, and second, to illustrate the invasion dynamics.
6.1 Parameter estimation and scaling analysis
A key prerequisite for the numerical simulations is a consistent estimation of the dimensionless parameters. Since precise quantitative measurements of many biological coefficients are not available, our goal is not to assign exact numerical values, but to identify biologically consistent scaling regimes that capture the dominant mechanisms of the system.
The model parameters arise from the interplay between different biological processes. By comparing characteristic timescales and transport mechanisms, we estimate their orders of magnitude and identify meaningful parameter ranges.
Time scales.
The reference timescale is given by the turnover rate of healthy tissue, $\mu_{H}^{-1}$. Experimental studies show that muscle injury triggers inflammatory responses that persist for several days to weeks, while signaling processes mediated by cytokines and chemokines occur on much shorter timescales [16, 38].
This leads to the hierarchy in which healthy tissue evolves on a timescale of weeks, macrophages act on timescales of days, and chemokines are produced and degraded over hours. Accordingly, the dimensionless rates satisfy
$\mu=\frac{\mu_{C}}{\mu_{H}}\sim O(10^{2}-10^{3}),\qquad\nu=\frac{\mu_{M}}{\mu_{H}}\sim O(10),\qquad\rho=\frac{r_{H}}{\mu_{H}}\sim O(1).$
Diffusion scales.
Transport processes in muscle tissue exhibit a clear separation between cellular migration and molecular diffusion. Immune cells migrate relatively slowly, while chemokines diffuse much faster through the extracellular environment. This yields
$D_{d}\sim 1,\qquad D_{m}\sim 5-20,\qquad D_{c}\sim 10^{2}-10^{3},$
which should be interpreted as effective diffusion coefficients at the tissue scale.
Space competition.
The parameter $\sigma=\frac{s}{K\mu_{H}}$ measures the intrinsic growth rate of healthy tissue relative to its carrying capacity. In physiological conditions, $\sigma$ is typically of order one or smaller, indicating that space limitation plays a relevant role in regulating tissue dynamics.
Damage signaling and chemotaxis.
The parameter $r=\frac{r_{C}K}{k_{c}\mu_{H}}$ quantifies the strength of chemokine production induced by damage. When $r\ll 1$, signaling is too weak to sustain immune recruitment and the system relaxes to the healthy state. Conversely, for $r\gg 1$, strong signaling promotes persistent immune activation and damage amplification. In the simulations, we fix $r=1$, corresponding to a balanced regime in which damage-induced chemokine production is neither negligible nor dominant. This choice allows us to focus on the interplay between damage amplification and repair without introducing an additional bias due to extreme signaling regimes.
The chemotactic sensitivity $\chi_{0}=\frac{\bar{\chi}k_{c}}{D_{H}}$ controls the spatial organization of the immune response: small values lead to diffuse inflammation, whereas larger values produce directed migration and sharper invasion fronts.
These considerations indicate that the transition from recovery to chronic degeneration is governed by the balance between damage amplification and tissue repair.
Interpretation of the computational domain.
The spatial variable is nondimensionalized using the diffusion length of healthy tissue,
$\ell_{H}=\sqrt{\frac{D_{H}}{\mu_{H}}},$
so that a dimensionless domain of size $L$ corresponds to a physical length
$L_{\mathrm{phys}}=L\,\ell_{H}.$
The computational domain should not be interpreted as the entire musculature of the body, but rather as an idealized portion of muscle tissue. In one spatial dimension this corresponds to a segment of tissue. For the numerical validation of the propagation speed, the domain must be sufficiently large compared with the intrinsic width of the invasion front. Large values of $L$ are therefore used to minimize boundary effects and accurately capture the asymptotic front dynamics.
Choice of parameter ranges.
Based on the above scaling arguments, we distinguish between different classes of parameters. The diffusion ratios and the timescale ratios are constrained by the biological hierarchy discussed above, whereas parameters such as $r$, $\chi_{0}$, $\kappa$, and $\alpha$ should be regarded as effective control parameters of the model. For these quantities, we do not aim at precise biological calibration, but rather at the identification of plausible exploratory ranges consistent with the qualitative mechanisms under investigation.
In particular, the values of $\mu$, $\nu$, $\rho$ and of the diffusion ratios reflect the separation between tissue turnover, immune-cell dynamics and chemokine signaling suggested by the biological literature [38]. By contrast, the ranges chosen for $r$, $\chi_{0}$, $\kappa$, and $\alpha$ are intended to explore different signaling and chemotactic regimes within a biologically plausible setting.
The parameter $c_{\epsilon}$ models a small basal inflammatory activity. In the simulations we fix $c_{\epsilon}=0.1$, consistently with the assumption that low-level inflammatory signaling may be present even in the absence of substantial damage, while strong immune recruitment is primarily induced by damage-dependent chemokine production.
The parameter
$\alpha=\frac{\bar{a}\,r_{M}K}{\mu_{H}\mu_{M}}$
measures the effective strength of immune-mediated damage. Unlike the timescale and diffusion ratios, $\alpha$ is not directly calibrated from biological data and should be interpreted as an effective control parameter.
For the reference parameter set used in the simulations, the invasion threshold given by (5.10) yields $\alpha_{c}\approx 31$. The simulations are therefore performed for values of $\alpha$ both below and above this threshold, typically in the range $\alpha\in[15,60]$, in order to capture the transition between decay and invasion.
We summarize below the parameter ranges adopted in the simulations (see Table 2).
Parameter Range Role
$D_{d}$ 0.510.5-1 effective tissue-scale diffusion
$D_{m}$ 5205-20 immune-cell motility
$D_{c}$ $10^{2}-10^{3}$ fast chemokine diffusion
$\delta$ 0.550.5-5 damage removal timescale
$\nu$ 5205-20 macrophage timescale ratio
$\mu$ $10^{2}-10^{3}$ chemokine timescale ratio
$\sigma$ $10^{-1}-1$ space competition
$\rho$ 0.130.1-3 repair timescale ratio
$r$ 0.1100.1-10 damage-induced signaling strength
$\chi_{0}$ 1101-10 chemotactic sensitivity
$\kappa$ 0.110.1-1 chemotactic saturation scale ratio
$\alpha$ 156015-60 immune-mediated damage intensity (explored around threshold)
$c_{\varepsilon}$ $0.1$ basal inflammatory activity
Table 2: Biologically motivated and exploratory ranges of the dimensionless parameters used in the numerical simulations. The first block contains parameters constrained by the timescale and transport hierarchy discussed in the text, the second block contains effective control parameters explored around the invasion threshold and $c_{\varepsilon}$ represents a fixed basal inflammatory activity.
6.2 Numerical validation of invasion dynamics
We validate the analytical results obtained from the linearized system by means of numerical simulations of the full nonlinear model.
We consider initial conditions consisting of small localized perturbations of the healthy equilibrium, corresponding to the early-stage regime in which invasion is initiated.
The simulations reveal a sharp transition between decay and invasion. When the threshold condition is not satisfied ($R_{d}<1$), perturbations decay and the system returns to the healthy equilibrium. Above the threshold, damage spreads across the domain in the form of a traveling front.
A quantitative comparison between analytical and numerical front speeds shows very good agreement over the parameter ranges explored here. This indicates that, in the early-stage regime, the invasion dynamics is well captured by the linearized system and that the front speed is accurately described by the linear spreading speed.
These results support the validity of the linear analysis for the onset of spatial spreading and confirm the relevance of the pulled-front picture in the early stage of disease progression.
6.2.1 Validation of the invasion threshold
We first validate the analytical prediction of the invasion threshold derived in Section 5. To this end, we consider initial conditions consisting of small localized perturbations of the healthy equilibrium.
Figure 2 shows the temporal evolution of the system for values of $\alpha$ below, near and above the invasion threshold.
Below the threshold, the initial perturbation decays and the system returns to the healthy equilibrium, confirming the stability of the healthy state with respect to small perturbations. Just above the threshold, we observe the onset of invasion: the damaged tissue grows and spreads across the domain, while the healthy tissue decreases. For larger values of $\alpha$, the invasion becomes faster and more pronounced.
These results confirm the existence of a sharp transition between decay and invasion, in agreement with the analytical threshold condition derived from the linearized system. The parameter values used in the simulations are reported in the caption of Figure 2.
Figure 2: Temporal and spatial dynamics of the system for different values of $\alpha$, illustrating the transition across the invasion threshold and the emergence of traveling-wave propagation. The three rows correspond, respectively, to the below-threshold ($\alpha=30$), near-threshold ($\alpha=32$) and strong invasion ($\alpha=50$) regimes. In the first row, the four panels (from left to right) represent the temporal evolution at $x=0$ of healthy tissue ($h$), damaged tissue ($d$), macrophages ($m$) and chemokines ($c$). In this regime, the perturbation decays and the system returns to the healthy equilibrium. In the second and third rows, the four panels show the spatial profiles of $h$, $d$, $m$, and $c$ at successive times, highlighting the formation and propagation of traveling waves connecting the healthy state to the diseased state. Near the threshold ($\alpha=32$), invasion is initiated but progresses slowly, while for larger values of $\alpha$ ($\alpha=50$) the invasion front propagates faster and becomes more pronounced. Simulations are performed with parameters $\sigma=1$, $\rho=0.9$, $c_{\epsilon}=0.1$, $\delta=1.1$, $\nu=7$, $k=r=1$, $\mu=168$, $D_{d}=1$, $D_{m}=10$, $D_{c}=1000$, and $\chi_{0}=5$. For this choice of parameters, the analytical threshold predicted by the linearized system is $\alpha_{c}\approx 31$. Initial conditions consist of a small localized perturbation of the healthy equilibrium.
6.2.2 Validation of the propagation speed
We now validate the analytical prediction of the propagation speed.
In the left panel of Figure 3, we compare the theoretical dispersion relation $s(\gamma)$ with numerical estimates of the front speed obtained from initial conditions of the form $e^{-\gamma x}$.
For values of $\gamma$ smaller than the critical decay rate $\gamma^{*}$ corresponding to the minimum of the dispersion curve, the numerically observed speed is larger than the linear prediction. In contrast, for $\gamma\geq\gamma^{*}$, the numerical speed converges to the minimal value $s^{*}$ predicted by the linear theory.
This behavior is characteristic of pulled front dynamics, where the asymptotic propagation speed is selected by the slowest decaying mode at the leading edge.
Figure 3: Left: Comparison between the analytical dispersion relation $s(\gamma)$ and numerical estimates of the front speed obtained from exponential initial conditions of the form $e^{-\gamma x}$. The minimal speed $s^{*}$ corresponds to the minimum of the dispersion curve and is selected for $\gamma\geq\gamma^{*}$. Right: Comparison between analytical and numerical front speeds for Gaussian initial conditions, for different domain sizes $L$. The agreement improves as $L$ increases, indicating that discrepancies observed for small domains are due to finite-size effects. The remaining parameters are as in Fig. 2, with $\alpha=32$ in the left panel.
The right panel of Figure 3 indicates that, for Gaussian initial conditions, the agreement between analytical and numerical speeds improves as the domain size $L$ increases. For sufficiently large domains, the numerical speed converges to the analytical prediction, confirming that deviations observed in smaller domains are due to finite-size effects.
We next examine the dependence of the propagation speed on the model parameters.
Figure 4: Comparison between the analytical wave speed predicted by the linearized system and the numerical front velocity. Left: propagation speed as a function of $\alpha$, with $\rho=0.9,\delta=1.1$. Increasing $\alpha$ leads to faster invasion. Middle: propagation speed as a function of $\rho$, with $\alpha=40,\delta=1.1$. Increasing $\rho$ slows down the propagation. Right: propagation speed as a function of $\delta$, with $\alpha=40,\rho=0.9$. Increasing $\delta$ slows down the propagation. The remaining parameters are set to $c_{\epsilon}=0.1$, $r=\kappa=\sigma=1$, $\nu=7$, $\mu=168$, $D_{d}=1$, $D_{m}=10$, $D_{c}=100$, and $\chi_{0}=5$, with domain size $L=800$.
An increase in $\alpha$, which quantifies the strength of immune-mediated damage, leads to a faster propagation of the damage front. Biologically, this corresponds to a more aggressive inflammatory response, which amplifies tissue degradation and accelerates disease spreading.
Conversely, increasing $\rho$, which represents the efficiency of damage-induced repair, reduces the propagation speed. This reflects the stabilizing role of regeneration mechanisms, which counteract the accumulation of damage and slow down its spatial expansion.
A similar effect is observed when increasing $\delta$, which corresponds to a faster removal of damaged tissue. Larger values of $\delta$ reduce the persistence of damage and therefore hinder its spatial propagation, leading to a decrease in the front speed. Overall, these results highlight the competition between damage amplification and damage removal (through repair and clearance) as the key mechanism controlling the rate of disease progression.
The agreement between analytical and numerical speeds remains remarkably accurate across the explored parameter range, showing that the dependence of the invasion speed on $\alpha$ and $\rho$ is already well captured by the linearized dynamics in the early stage of disease progression.
7 Discussion and Conclusions
In this work, we developed and analyzed a spatially extended mathematical model for DMD driven by immune-mediated tissue damage. Starting from existing ODE-based formulations, we introduced a reaction system incorporating tissue capacity constraints, damage-induced regeneration and saturating immune activation. These mechanisms capture essential features of muscle repair and inflammatory regulation and provide a consistent basis for spatial modeling.
After extending the model to a reaction–diffusion–chemotaxis framework, we established fundamental analytical properties, including positivity, boundedness and positive invariance of a biologically admissible domain. These results ensure the well-posedness of the system and the physical relevance of its solutions. We also characterized the steady states of the nondimensionalized model, identifying biologically meaningful equilibria corresponding to healthy and pathological conditions.
A central aspect of the present study is the analysis of early-stage disease dynamics through linearization around the healthy equilibrium. Our results show that diffusion does not induce Turing instabilities, indicating that spatial heterogeneity cannot arise from classical diffusion-driven mechanisms. Instead, pathological progression occurs through invasion processes, in which localized damage spreads across the tissue. This suggests that the heterogeneous patterns observed in imaging studies are not generated by intrinsic pattern-forming mechanisms [31, 35], but rather by the spatial expansion of localized regions of damage.
We derived explicit conditions for the onset of invasion and characterized the minimal propagation speed of pathological fronts. These results provide a quantitative description of early disease spreading and show that the invasion process is governed by a pulled-front mechanism controlled by the linearized dynamics. The existence of an invasion threshold provides a possible mechanistic interpretation of the transition from compensated muscle function to progressive degeneration observed clinically. Numerical simulations of the full nonlinear system support these findings and confirm the predicted transition between decay and invasion, as well as the dependence of the propagation speed on key model parameters. In particular, the dependence of the front speed on parameters associated with immune-mediated damage and tissue repair suggests a direct link between inflammatory activity and the rate of disease progression.
The present work also highlights several limitations and open problems. The existence and stability properties of the diseased equilibrium remain difficult to characterize in full generality, due to the nonlinear coupling and the large number of parameters. Nevertheless, in a parameter regime consistent with the early-stage scenario considered here, the pathological equilibrium can be partially characterized and its existence is directly linked to the invasion threshold. A more detailed bifurcation analysis with respect to biologically relevant parameters could provide deeper insight into long-term disease outcomes. Extending the analysis beyond the early-stage regime, particularly in the vicinity of pathological equilibria, is a natural next step. Numerical evidence indicates the presence of bistable regimes in which healthy and pathological states coexist, and a deeper examination of the associated stability and bifurcation structure could further clarify the mechanisms governing long-term disease progression.
The role of chemotaxis also deserves further consideration. Although it does not significantly influence the early-stage invasion dynamics analyzed here, its inclusion provides a structurally consistent framework for exploring regimes with stronger inflammatory activity. In such regimes, directed immune migration may become more relevant and contribute to the emergence of spatially heterogeneous patterns, potentially related to localized infiltration and fatty replacement observed in magnetic resonance imaging of DMD patients [9]. More generally, the present results indicate that early-stage spatial degeneration is primarily driven by local damage–inflammation feedback, rather than by directed immune cell migration.
While diffusion does not induce Turing instabilities around the healthy equilibrium, this does not exclude the possibility that diffusion-driven pattern formation may arise in the vicinity of pathological states. In regimes characterized by sustained inflammation and higher levels of damage, the interplay between nonlinear reaction terms, diffusion and chemotaxis may lead to the emergence of spatially heterogeneous structures. Such patterns could provide a possible mechanistic explanation for the localized infiltration and fatty replacement observed in magnetic resonance imaging of DMD patients. Identifying the corresponding parameter ranges and dynamical regimes remains an open problem.
Moreover, the present model does not explicitly account for spatial heterogeneity in tissue properties, vascularization or mechanical constraints, which are known to influence muscle degeneration. Incorporating these effects may lead to more realistic descriptions of disease progression.
From a biological perspective, the model provides a simplified representation of complex immune and repair processes. Extensions including additional immune cell populations, signaling pathways or feedback mechanisms may improve biological realism, at the cost of increased mathematical complexity.
Finally, the proposed framework can be adapted to other chronic inflammatory or degenerative diseases characterized by damage-driven immune responses, providing a flexible basis for investigating spatial disease dynamics beyond DMD.
In conclusion, we introduced a mathematically tractable spatial model capturing key features of immune-mediated muscle degeneration. By combining analytical techniques and numerical simulations, we provided new insight into early disease spreading and invasion phenomena, highlighting the role of damage-driven mechanisms in shaping spatial disease dynamics. These results offer a quantitative framework for understanding the onset of spatial degeneration and clarify the mechanisms underlying early-stage disease progression.
Acknowledgements
This work has been partially supported by the European Union – Next Generation EU: PRIN project MUR PrinPNRR 2022, project code P202254HT8, Grant No. CUP B53D23027760001; by the PRIN project MUR PrinPNRR 2022, project code P2022Z7ZAJ, Grant No. CUP B53D23027930001 and by the project the PRIN project MIUR Prin 2022, project code 1074 2022M9BKBC, Grant No. CUP B53D23009350006. The authors also gratefully acknowledge the financial support of GNFM-INdAM.